### v1

# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

Abstract
Recent advances in data-driven evolutionary algorithms (EAs) have demonstrated the potential of leveraging data to improve optimization accuracy and adaptability. Nevertheless, most existing approaches remain dependent on handcrafted heuristics, which limits their generality and automation. To address this challenge, we propose Evolutionary Generative Optimization (EvoGO), a fully data-driven framework empowered by generative learning. EvoGO streamlines the evolutionary optimization process into three stages: data preparation, model training, and population generation. The data preparation stage constructs a pairwise dataset to enrich training diversity without incurring additional evaluation costs. During model training, a tailored generative model learns to transform inferior solutions into superior ones. In the population generation stage, EvoGO replaces traditional reproduction operators with a scalable and parallelizable generative mechanism. Extensive experiments on numerical benchmarks, classical control problems, and high-dimensional robotic tasks demonstrate that EvoGO consistently converges within merely 10 generations and significantly outperforms a wide spectrum of optimization approaches, including traditional EAs, Bayesian optimization, and reinforcement learning based methods. Source code will be made publicly available.
Index Terms:
Generative learning, learning-based evolutionary algorithm, large-scale expensive optimization
I Introduction
Many real-world optimization tasks involve objectives that are accessible only through simulations, legacy software, or physical experiments [1, 2]. Such complex problems are typically non-analytical and lack gradient information, presenting significant challenges for traditional optimization techniques. Evolutionary Algorithms (EAs) have emerged as a versatile and widely adopted framework for tackling such problems, with demonstrated success in domains such as hyperparameter tuning, biomedical engineering, and aerospace design [3]. Classical EAs maintain a population of candidate solutions and apply handcrafted variation operators (e.g., crossover and mutation) to iteratively explore the search space [4]. Although effective in many scenarios, these operators often struggle in high-dimensional or rugged landscapes, such as those found in complex robotic control tasks [5].
To enhance the performance of conventional EAs, researchers have introduced domain-specific heuristics into them. For instance, Evolutionary Strategies (ES)[6] and Estimation of Distribution Algorithms (EDAs)[7] employ statistical modeling to guide the search process. Despite their success in specialized contexts, these methods still rely on manually crafted rules and assumptions, thereby limiting their scalability, generality, and adaptability across diverse problem domains. With the increasing scale and complexity of modern engineering systems, along with the growing availability of data, conventional EAs face mounting challenges in navigating high-dimensional search spaces and addressing large-scale optimization tasks. This has led to a rising demand for more powerful EAs capable of handling increasingly sophisticated problems. To meet this demand, researchers have explored the integration of machine learning techniques into evolutionary computation, giving rise to a new class of learnable EAs [8]. These methods are generally categorized into two groups: non-data-driven and data-driven ones. The non-data-driven methods  [9, 10, 11, 12] typically enhance algorithmic components through predefined rules or online adaptation, but without systematically leveraging the data collected during the optimization process. In contrast, data-driven methods  [13, 14, 15, 16, 17] explicitly model and utilize the historical search data (e.g., evaluated solutions and their fitness values) to inform and guide future search decisions in a more principled and efficient manner.
A notable subclass within the data-driven category is generative modeling-based evolutionary algorithms (GMEAs) [18, 19, 20], which have recently attracted increasing attention due to the promising capabilities of generative models [21, 22, 23, 24]. By coupling the exploratory nature of evolutionary search with the expressive power of generative modeling, GMEAs establish a positive feedback loop between solution generation and model learning. This synergy offers notable potential for solving complex and high-dimensional optimization tasks [4]. However, existing GMEAs are still in their infancy and face three critical challenges: (i) The integration of generative models into evolutionary workflows is often ad hoc and lacks a principled mechanism, thereby necessitating the use of additional heuristic modules. (ii) The training objectives of generative models are frequently borrowed from unrelated domains such as image or text generation, thereby resulting in misalignment with the goals of evolutionary optimization. (iii) Most current methods rely solely on limited population data, which constrains training stability and generalization, particularly in expensive optimization scenarios where evaluation budgets are tight.
These challenges suggest that a fully data-driven evolutionary optimization framework (i.e., one that replaces handcrafted heuristics with end-to-end learning-driven mechanisms) remains an open research goal. In response, we propose Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework that eliminates reliance on traditional heuristic operators through generative learning. Our key contributions are summarized as follows:
  • We design a paired generative architecture in which forward and inverse models collaborate to produce high-quality candidate solutions. This setup enables direct mapping from historical populations to new individuals, entirely replacing traditional crossover and mutation heuristics.
  • We introduce a tailored loss function that combines reconstruction accuracy, distributional alignment, and directional guidance. This composite objective explicitly steers the generative process toward promising regions of the search space.
  • We develop a learning-based sample synthesis strategy that generates informative training data from limited fitness evaluations. This mechanism supports stable model training even under small population sizes and tight evaluation budgets.
The remainder of the paper is organized as follows. Section II reviews existing approaches for complex optimization and discusses their relationships to EvoGO. Section III introduces the proposed model architecture, loss function, and learning strategies. Section IV presents a comprehensive empirical evaluation, including performance comparisons and ablation studies. Finally, Section V concludes the paper.
II Related Work
II-A Non-Data-Driven Approaches
EAs are among the most common non-data-driven approaches for complex optimization. They operate without leveraging data-driven models or learning mechanisms, relying instead on fixed heuristics to explore the search space. These methods can be broadly classified into two categories: (i) those that generate new solutions through handcrafted search operators, and (ii) those that probabilistically model the distribution of promising solutions to guide sampling.
The first category encompasses classic algorithms such as genetic algorithms [9], genetic programming [10], differential evolution [25], and particle swarm optimization [26]. These methods generate new candidate solutions using handcrafted operators like crossover, mutation, or velocity updates. While widely adopted, they rely heavily on domain knowledge and manual tuning, which can limit their flexibility and scalability. To address some of these limitations, the second category i.e, probabilistic modeling-based methods, attempts to guide the search by learning distributions over promising solutions. This category includes two representative families: Evolutionary Strategies (ES) and Estimation of Distribution Algorithms (EDAs).
ES methods emerged from early work on stochastic optimization [11] and have since evolved through several key innovations. Among them, the Covariance Matrix Adaptation ES (CMA-ES) introduced an adaptive mutation mechanism that adjusts the search distribution based on population statistics [27], and was later extended to a fully derandomized formulation [28]. More recently, ES has been applied to large-scale machine learning problems, such as in OpenAI-ES [29], which demonstrated impressive scalability and parallelism in reinforcement learning settings.
In parallel, EDAs integrate statistical modeling into evolutionary search by explicitly learning the distribution of high-quality solutions. The seminal Population-Based Incremental Learning (PBIL) algorithm [30] introduced this idea in discrete spaces, followed by the Univariate Marginal Distribution Algorithm (UMDA) [31], which simplifies modeling by assuming independence among variables. Further developments sought to capture dependencies through more expressive models, such as the extended compact genetic algorithm [32] and the linkage tree genetic algorithm [33].
Despite their conceptual elegance, both ES and EDA methods remain fundamentally non-data-driven in nature. Their update and sampling procedures are still crafted manually and follow rigid algorithmic rules. This restricts their adaptability in diverse or complex scenarios [6, 7]. To enhance flexibility and adaptability, some studies have explored incorporating reinforcement learning (RL) into EAs, aiming to automate the selection or configuration of evolutionary operators [34, 35, 36, 37, 38]. These methods typically employ RL agents to dynamically choose operators or strategies based on past performance. While such integration introduces a level of adaptivity, the learning is still external to the optimization process itself, and often comes with substantial computational overhead and poor sample efficiency [39]. Thus, they do not fundamentally depart from the non-data-driven paradigm, and the search dynamics remain largely heuristic in nature.
II-B Data-Driven Apporoaches
While EAs primarily rely on handcrafted heuristics and static update rules, recent advances have increasingly turned to data-driven approaches that aim to learn from the information accumulated during the optimization process. These approaches leverage machine learning models to improve the efficiency, scalability, or adaptability of evolutionary optimization. Among them, two major categories have emerged: Bayesian Optimization, which constructs surrogate models to guide the search via probabilistic inference, and learnable EAs, which integrate machine learning modules into the evolutionary cycle itself.
1) Bayesian Optimization
Bayesian Optimization (BO) traditionally employs Gaussian processes to model the objective function and constructs acquisition functions to determine promising query points [40]. To improve performance in large-scale and high-dimensional settings, substantial efforts have been made to enhance the scalability of BO [41, 42, 43, 44, 45, 46, 47, 48]. Despite these advances, significant computational overhead persists when BO is applied to complex tasks on modern high-throughput computing platforms.
To address this, researchers have proposed multi-point acquisition strategies to enable batch querying [49, 50, 51, 52]. Among these, Thompson Sampling (TS) [53] has gained popularity for parallel BO [54, 55, 56, 57]. For example, Eriksson et al. integrated TS with a trust region method to achieve scalable BO for both constrained and unconstrained problems, supporting batch sizes of up to 100. Nonetheless, in most cases, the batch sizes remain relatively modest, resulting in underutilization of modern computational resources.
In parallel, Artificial Neural Networks (ANNs) have been explored to improve the batch efficiency of BO [58, 59, 60]. These approaches, however, often repurpose the surrogate model as a binary classifier rather than directly generating new candidate solutions. For instance, Tiao et al. and Oliveira et al. proposed density-ratio-based methods, while Swersky et al. adopted reinforcement learning to train a selection policy for discrete domains. Despite their novelty, these methods still fundamentally rely on direct pointwise optimization for generating new queries [61], which hinders the full scalability of BO.
2) Learnable EAs
Recent advances have witnessed a growing interest in incorporating machine learning techniques into evolutionary algorithms, giving rise to the paradigm of learnable EAs. These methods can be broadly categorized into two main classes: EAs with learnable evaluators, which utilize predictive models to approximate fitness values; and EAs with learnable generators, which aim to improve offspring generation through learned transformations [4].
EAs with Learnable Evaluators
These methods are also known as the Surrogate-Assisted EAs (SAEAs), which are particularly suited for expensive optimization problems where fitness evaluations involve costly simulations. Early frameworks in model-based evolutionary computation [62, 63] established the integration of approximation models to reduce evaluation cost. Traditional SAEAs often employ simple surrogate models such as Radial Basis Functions (RBFs) and Support Vector Machines (SVMs) [64], which are limited in capturing the complexity of high-dimensional landscapes. The introduction of Gaussian Processes (GPs) addressed some of these limitations by enabling uncertainty-aware acquisition functions. For example, Liu et al. [65] combined GPs with EA operators to balance exploration and exploitation. More recent hybrid methods show improved scalability. SAMSO [66] incorporates dual swarm populations with a surrogate model to handle up to 200 decision variables, while AutoSAEA [67] dynamically selects among multiple surrogate models for robust performance across benchmark and real-world tasks. Nonetheless, existing SAEAs often rely heavily on surrogate predictions for candidate selection, which can be computationally intensive and susceptible to model bias [68].
EAs with Learnable Generators
For simplicity, we refer to these methods as Generative Modeling-based Evolutionary Algorithms (GMEAs), which aim to enhance evolutionary optimization by replacing handcrafted variation operators with learned generative mechanisms. These methods have demonstrated promising results across a range of single- and multi-objective problems. Current research in this area predominantly follows two directions: (i) leveraging unsupervised learning for variable grouping and dimensionality reduction [69, 70, 18], and (ii) employing supervised or reinforcement learning to guide the directional improvement of candidate solutions [71, 72, 73, 19, 20]. The first line of work trains generative models to discover latent structure within the search space, often to facilitate more efficient variation. For example, Pan et al. [18] proposed a manifold-learning-based approach that restricts mating operations to low-dimensional subspaces, thereby improving optimization efficiency. The second line focuses on guiding evolutionary search using data-driven feedback. He et al. [73] introduced a GAN-based multi-objective EA where a discriminator distinguishes between real and synthetic populations to train a generator that produces new candidates. Similarly, Liu et al. [19] employed reinforcement learning to adaptively select variation operators based on individual context and historical performance. While GMEAs offer significant potential, they also present several open challenges. These include the difficulty of seamlessly integrating generative models into the evolutionary loop, ambiguities in defining effective training objectives, and a dependence on large population sizes to ensure model generalization and stability.
II-C Discussion
Recent advances in learnable EAs have demonstrated the promise of integrating machine learning into population-based optimization. However, most existing approaches adopt a modular design, treating key components such as fitness prediction, offspring generation, and selection as independent units. While this separation simplifies implementation, it inevitably leads to fragmented learning dynamics and underutilization of the information shared across modules.
In principle, these components are inherently interdependent and can benefit from tighter integration. For example, a shared latent space could act as a unified representation that bridges evaluators and generators: the surrogate model refines this space by encoding fitness-related feedback, while the generator explores it to produce high-potential solutions. Such a design encourages the emergence of a self-reinforcing feedback loop, enabling evaluation and generation to co-adapt based on shared experience, rather than evolving in isolation. Despite this potential, current methods still rely heavily on manually crafted heuristics, rigid coordination rules, and loss functions inherited from unrelated domains (e.g., image synthesis or language modeling). This not only introduces mismatches between local learning objectives and the global goal of optimization, but also undermines the transferability and robustness of the resulting models across different problem settings.
To overcome existing limitations, a promising approach is to develop an end-to-end learning framework that integrates representation learning, solution generation, and fitness evaluation within a cohesive pipeline. Such a framework would ideally (i) eliminate reliance on handcrafted coordination mechanisms, (ii) ensure alignment between the learning objectives of each module and the global optimization goal, and (iii) fully leverage all available information to improve data efficiency and adaptability.
III Method
This section introduces the overall framework of EvoGO, followed by detailed descriptions of each key component.
Figure 1: The overall framework of EvoGO. The historical dataset $\mathcal{D}$ is composed of $\{\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}\}$, where each $\mathcal{D}_{t}=\left\{(\bm{x}_{t,i},y_{t,i})\mid 1\leq i\leq N\right\}$ , with $\bm{x}_{t,i}$ and $y_{t,i}$ denoting the population and its fitness, respectively. In data preparation phase, the historical population is first selected to form the training set of current generation and augmented by an optional learning-based data augmentation strategy. Then, it is separated into the superior and inferior groups with smaller and larger fitnesses, respectively, before a Cartesian product is applied to generate the paired training dataset. In model training phase, the framework first trains a surrogate model solely for landscape estimation. Then, the surrogate model is frozen and combined with a pair of generative models to form a composite model for effective generative model training. Later, the composite model is trained via the proposed training loss tailored for comoplex optimization. In population generation phase, the inferior-to-superior generative model is used to generate the new solutions in parallel.
III-A Overall Framework
Algorithm 1 Data Preparation
1:Input: historical data $\mathcal{D}$, dataset split factor $\eta\in(0,1)$, sliding window factor $\epsilon>=0$,
Input: data augmentation threshold $N$ and factor $\rho$
2:Output: $\mathcal{D}_{\text{SM}}$ for surrogate model training, $\mathcal{D}_{\text{GM}}$ for generative model training
3: /************** Elite Selection **************/
[…]
6: /************** Data Augmentation (Optional) **************/
7:if $|\mathcal{D}_{0}|<N$ then
8:  $\tilde{f}\leftarrow$ trained surrogate model on $\mathcal{D}_{\text{SM}}$
9:  $\bm{\mathcal{X}}^{\prime}\leftarrow$ randomly generate $\rho\cdot|\mathcal{D}^{\prime}|$ solutions according to the distribution of solutions in $\mathcal{D}^{\prime}$
10:  $\mathcal{D}^{\prime}\leftarrow\{(\bm{x},\mu_{t}(\bm{x}))|\bm{x}\in\bm{\mathcal{X}}\}$ where $\mu_{t}(\bm{x})=\mathbb{E}[\tilde{f}(\bm{x})]$
11:end if
12: /************** Paired Set Construction **************/
[…]
As illustrated in Fig. 1, EvoGO comprises three key phases: data preparation, model training, and population generation. To mitigate the risk of overfitting caused by limited population data, the data preparation phase incorporates a pairwise data augmentation strategy, enhancing training diversity without introducing additional evaluation costs. In the model training phase, a composite model architecture and a tailored loss function are employed to align the generative training objectives with the goals of evolutionary optimization, thereby addressing the objective misalignment issue. Moreover, to eliminate reliance on handcrafted rules, EvoGO adopts an end-to-end generative paradigm, replacing the heuristic components commonly used in traditional GMEAs. Within this data-driven framework, data preparation replaces conventional environmental selection; model training substitutes for basic surrogate modeling; and population generation takes the place of heuristic-based reproduction.
III-B Data Preparation
The data preparation phase in EvoGO aims to construct high-quality and structured training data for subsequent model training. Specifically, it (i) selects elite samples from historical populations to ensure surrogate accuracy, (ii) optionally augments the data to improve diversity under limited evaluations, and (iii) generates a paired dataset to provide directional learning signals for the generative model.
As detailed in Algorithm 1, the data preparation process comprises the following three steps:
  • Elite Selection: A high-quality dataset $\mathcal{D}_{\text{SM}}$ is constructed by selecting individuals with the highest fitness values from all historical populations $\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}$. This ensures accurate surrogate modeling by concentrating on representative and promising solutions.
  • Data Augmentation (Optional): To mitigate the limited data availability in expensive optimization scenarios, an optional learning-based data augmentation strategy is employed. Specially, a Variational AutoEncoder (VAE) is used to generate diverse synthetic samples that resemble the current population, thereby expanding the generative model dataset $\mathcal{D}_{\text{GM}}$ when the original population is insufficient.
  • Paired Set Construction: Following augmentation (if applied), the surrogate model estimates the fitness of all candidate solutions. The samples are then ranked and partitioned into superior ($\mathcal{D}_{\text{GM}^{+}}$) and inferior ($\mathcal{D}_{\text{GM}^{-}}$) subsets based on a predefined split ratio $\eta$. Their Cartesian product forms the paired training set $\mathcal{D}_{\text{GM}}$, enabling the composite model to learn both effective descent directions and adaptive step sizes.
III-C Model Training
In the model training phase, we develop a composite model architecture that integrates a surrogate model with a pair of generative models. To support this structure, we define a dedicated composite loss function that enables directional learning and optimization-aware training. This phase is designed as an online process, where lightweight training is conducted during each evolutionary generation. Although the training involves three model components, only the generative model $\bm{\gamma}$ is used during population generation. The overall training procedure is described below, followed by detailed explanations of the model architecture and loss formulation.
1) Model Architecture
Conventional GMEAs typically rely on generative models in conjunction with heuristic procedures, often neglecting the optimization guidance available from surrogate-estimated objective landscapes. To address these limitations, we propose a composite model architecture (see Fig. 2) that integrates a surrogate model $\tilde{f}$ with a forward generative model $\bm{\gamma}$, which maps inferior solutions to superior ones, and a pseudo-inverse model $\bm{\gamma}^{\prime}$ that performs the reverse mapping. In contrast to GANs [74] and CycleGANs [75], which rely on discriminators, our approach leverages the surrogate model to impose optimization-aware learning via gradient back-propagation.
Given a population of solutions, our model is designed to generate new solutions that potentially improve function values. The dual generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ implement the generative loss $\mathcal{L}_{\text{gen}}$ (Eq. 2 and Eq. 7), while the surrogate model $\tilde{f}$ addresses the optimization loss $\mathcal{L}_{\text{opt}}$. Initially, the surrogate model is established and remains constant throughout the training process; $\bm{\gamma}$ processes inferior solutions to generate estimated superior ones, and $\bm{\gamma}^{\prime}$ operates in reverse, both facilitating the computation of the similarity loss $\mathcal{L}_{\text{sim}}$. Subsequently, the solutions generated by $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are reconstructed by cycling back to each other, facilitating the computation of reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5). Additionally, the generated solutions ($\bm{\gamma}(\bm{p})$ and $\bm{\gamma}^{\prime}(\bm{q})$) may also contribute to the optimization loss $\mathcal{L}_{\text{opt}}$ when evaluated against the surrogate. This configuration facilitates the transfer of optimization landscape estimation from the surrogate to the generative models via back-propagation, enabling $\bm{\gamma}$ to autonomously generate optimized populations.
In summary, the architecture integrates a surrogate model $\tilde{f}$, alongside a generative model $\bm{\gamma}$ and its pseudo-inverse $\bm{\gamma}^{\prime}$. $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are specifically trained for generating and reconstructing the superior and inferior solutions, respectively, while $\tilde{f}$ focuses on the estimation of the optimization landscape. Together, their synergistic training significantly enhances $\bm{\gamma}$’s ability to generate optimized solutions.
2) Loss Function
As illustrated in Fig. 2, we define a composite loss function that integrates two principal components: the generative loss $\mathcal{L}_{\text{gen}}$ and the optimization loss $\mathcal{L}_{\text{opt}}$:
$\mathcal{L}=\mathcal{L}_{\text{gen}}+\lambda\mathcal{L}_{\text{opt}},$ (1)
where $\lambda$ is a trade-off parameter that balances generative learning and optimization guidance.
The generative loss $\mathcal{L}_{\text{gen}}$ comprises two components: a similarity loss $\mathcal{L}_{\text{sim}}$, which encourages consistent latent representations between paired samples, and a reconstruction loss $\mathcal{L}_{\text{rec}}$, which enforces structural consistency in the generated outputs:
$\mathcal{L}_{\text{gen}}=\mathcal{L}_{\text{sim}}+\lambda_{1}\mathcal{L}_{\text{rec}}.$ (2)
The optimization loss $\mathcal{L}_{\text{opt}}$ consists of an expectation loss $\mathcal{L}_{\text{mean}}$, which penalizes poor average fitness among generated samples, and a variance loss $\mathcal{L}_{\text{std}}$, which promotes diversity:
$\mathcal{L}_{\text{opt}}=\mathcal{L}_{\text{mean}}+\lambda_{2}\mathcal{L}_{\text{std}},$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the contributions of the reconstruction and variance terms, respectively.
In generative learning, the similarity loss $\mathcal{L}_{\text{sim}}$ is widely employed to ensure that generated outputs closely resemble target samples [76, 75, 77, 78]. For our generative model $\bm{\gamma}$, which transforms a source solution $\bm{p}$ to a target solution $\bm{q}$, we adopt a mean squared loss:
$\mathcal{L}_{\text{sim}}=\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}.$ (4)
This aligns with conventional generative paradigms, where models learn mappings from latent or random variables to target domains [74, 79, 80]. However, when applied to optimization problems, such models often struggle with the intrinsic trade-off between exploration and exploitation, especially within the EA framework [4, 81, 82].
Alternatively, domain adaptation models, which transform solutions between source and target domains [83, 84, 75, 85], have shown promising results in certain optimization scenarios [82, 86]. Nonetheless, without reconstruction constraints, these models are susceptible to mode collapse [76, 77, 78], compromising training stability. To mitigate this issue, we introduce a paired generative model $\bm{\gamma}^{\prime}$ that serves as a pseudo-inverse of $\bm{\gamma}$. This yields a reconstruction loss:
$\mathcal{L}_{\text{rec}}=\|\bm{\gamma}^{\prime}(\bm{\gamma}(\bm{p}))-\bm{p}\|^{2}+\|\bm{\gamma}(\bm{\gamma}^{\prime}(\bm{q}))-\bm{q}\|^{2},$ (5)
thereby leading to the similarity loss symmetrized as:
$\mathcal{L}_{\text{sim}}=\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}+\|\bm{\gamma}^{\prime}(\bm{q})-\bm{p}\|^{2}.$ (6)
The generative loss $\mathcal{L}_{\text{gen}}$, particularly the reconstruction term, plays a pivotal role in stabilizing training and preserving semantic fidelity. This is crucial in optimization contexts, where the model must accurately learn transformations from inferior to superior regions of the search space without drifting off-manifold. To ensure this, the weight $\lambda_{1}$ associated with the reconstruction loss is set to dominate the loss function, emphasizing faithful reconstruction over mere proximity.
By integrating the similarity and reconstruction losses for the bidirectional models $\{\bm{\gamma},\bm{\gamma}^{\prime}\}$, we obtain the full generative loss:
$\mathcal{L}_{\text{gen}}$ $=\left(\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}+\|\bm{\gamma}^{\prime}(\bm{q})-\bm{p}\|^{2}\right)$
$\quad+\lambda_{1}\left(\|\bm{\gamma}^{\prime}(\bm{\gamma}(\bm{p}))-\bm{p}\|^{2}+\|\bm{\gamma}(\bm{\gamma}^{\prime}(\bm{q}))-\bm{q}\|^{2}\right).$ (7)
This formulation adheres to the principles of domain-adaptive generative learning, where $\bm{p}$ and $\bm{q}$ are samples from source (inferior) and target (superior) domains, respectively. The dataset construction process is detailed in Section III-B, and an alternative interpretation of Eq. 7 from a probabilistic (Gaussian) perspective is provided in Supplementary Document A.
Incorporating our generative model into the EA framework requires transforming traditional optimization objectives into suitable loss function terms for complete data-driven process. To some extent, this is similar to the infill criteria in the Surrogate-Assisted Evolutionary Algorithms (SAEAs), where the surrogate model is used to select the most promising solution(s) from a population of candidates [68]. Such strategies have been widely studied in the literature, and one of the most common approach is to use the Upper/Lower Confidence Bound (UCB/LCB) [41] as the infill criteria [65]:
$\alpha(\bm{p})=\mathbb{E}\left[\tilde{f}(\bm{p})\middle|\bm{p}\right]+\lambda_{2}\mathbb{V}\left[\tilde{f}(\bm{p})\middle|\bm{p}\right]^{1/2},$ (8)
where $\alpha$ is the UCB/LCB strategy to be applied to each candidate $\bm{p}$, $\tilde{f}$ is the surrogate of the objective function which we are trying to minimize and $\lambda_{2}$ is a hyperparameter as mentioned in Eq. 3. In traditional SAEAs, all candidate solutions are evaluated against the Eq. 8 to find the one(s) with the minimal value(s) of $\alpha(\bm{p})$, which is then evaluated by the true objective function.
Figure 2: Overview of the model architecture and composite loss functions $\mathcal{L}$ in EvoGO. The top, middle, and bottom rows illustrate the computational workflows for the similarity loss $\mathcal{L}_{\text{sim}}$, the reconstruction loss $\mathcal{L}_{\text{rec}}$, and the optimization loss $\mathcal{L}_{\text{opt}}$, respectively. The input pair $(\bm{p},\bm{q})$ corresponds to solutions with inferior and superior fitness, respectively.
However, the conventional use of acquisition functions such as UCB or LCB is not directly applicable to our generative model $\bm{\gamma}$. This is because the acquisition strategy must be adapted to incorporate both the surrogate model and the generative model. Inspired by principles from meta-learning [87, 88], we propose to reinterpret the acquisition criterion as a training loss for the generative model. Without loss of generality, we define the generative model as $\bm{\gamma}(\cdot;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, parameterized by $\bm{\phi}$, which generates a new solution $\bm{x}_{t+1}$ from an existing input $\bm{x}$:
$\bm{x}_{t+1}=\bm{\gamma}\left(\bm{x};\operatorname{argmin}_{\bm{\phi}}\mathcal{L}(\bm{\phi})\right),$
$\text{where}\quad\mathcal{L}(\bm{\phi})=\alpha(\bm{\gamma}(\bm{x};\bm{\phi})).$ (9)
In this formulation, the acquisition function $\alpha(\cdot)$ is used to define the optimization objective for training the parameters $\bm{\phi}$ of the generative model. Through this transformation, the task of maximizing the acquisition function is seamlessly integrated into the learning process of the generative model, allowing it to internalize exploration-exploitation trade-offs during training.
Thus, by replacing the objective function $f$ with it surrogate model $\tilde{f}$, we derive the optimization loss of UCB/LCB as:
$\mathcal{L}_{\text{opt}}=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2},$ (10)
where $\mathbb{E}$ and $\mathbb{D}$ represent the mean and variance of the surrogate model’s output, respectively. This loss ties the generative model’s parameters directly to the surrogate’s predictions, facilitating a robust integration of surrogate model into a generative modeling context. However, since the transformation of the infill criteria to the optimization loss is not guaranteed to be lossless, the inherent balance between exploration and exploitation may be compromised. Hence, we introduce an explicit variance optimization loss composed of UCB/LCB and Knowledge Gradient (KG) [89]:
$\mathcal{L}_{\text{opt}}$ $=\alpha_{\text{KG}}(\bm{p})$
$=\mathbb{E}\left[\mu^{*}_{t+1}-\mu^{*}_{t}\mid\bm{p}\right]+\lambda_{2}\mathbb{D}\left[\mu^{*}_{t+1}-\mu^{*}_{t}\mid\bm{p}\right]^{1/2}$
$=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2}+$
$\lambda_{2}\mathbb{C}\left[\tilde{f}_{t}(\bm{\gamma}^{\prime}\circ\bm{\gamma}(\bm{p})),\tilde{f}_{t}(\bm{q})\right]+\lambda_{2}\mathbb{C}\left[\tilde{f}_{t}(\bm{\gamma}\circ\bm{\gamma}^{\prime}(\bm{q})),\tilde{f}_{t}(\bm{p})\right],$ (11)
where $\mathbb{C}$ is the correlation function. Since the loss function is to be minimized, the existence of correlation function requests that the parameter $\lambda_{2}$ is negative, which is not necessarily the case in Eq. 10. Supplementary Document A will elaborate the details of the justification of Eq. 11.
It is important to note that, unlike in conventional SAEAs, the surrogate model $\tilde{f}$ in our framework is not directly used to evaluate the quality of generated solutions. Instead, it serves as a guidance mechanism during the training of the generative model. Furthermore, given the wide variety of infill criteria proposed in the literature, the acquisition function in Eq. 8 can be flexibly substituted with alternative strategies. This substitution naturally leads to different formulations of the optimization loss $\mathcal{L}_{\text{opt}}$, thereby allowing our framework to accommodate various exploration-exploitation trade-offs.
3) Training Procedure
The training procedure is summarized in Algorithm 2. It begins with training the surrogate model $\tilde{f}$ on the elite dataset $\mathcal{D}_{\text{SM}}$, after which the surrogate is frozen to ensure stable guidance throughout the optimization process. The paired generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are subsequently optimized using mini-batches sampled from the paired dataset $\mathcal{D}_{\text{GM}}$, which is constructed during the data preparation phase (Algorithm 1).
Algorithm 2 Model Training
1:Input: maximum training epochs $E$, optimizer opt, training datasets $\mathcal{D}_{\text{SM}}$ and $\mathcal{D}_{\text{GM}}$
2:Output: trained generative model $\bm{\gamma}$
[…]
4:for $e=1$ to $E$ do
5:  for each pair $((\bm{x}_{-},y_{-}),(\bm{x}_{+},y_{+}))$ in $\mathcal{D}_{\text{GM}}$ do
6:   $(\bm{p},y_{\bm{p}}),(\bm{q},y_{\bm{q}})\leftarrow(\bm{x}_{-},y_{-}),(\bm{x}_{+},y_{+})$
7:   $\mathcal{L}_{b}\leftarrow\mathcal{L}(\bm{\phi},\bm{\phi}^{\prime})$ where $\mathcal{L}$ is Eq. 1
8:   $\bm{\phi},\bm{\phi}^{\prime}\leftarrow\texttt{opt}(\bm{\phi},\nabla_{\bm{\phi}}\mathcal{L}_{b}),\texttt{opt}(\bm{\phi}^{\prime},\nabla_{\bm{\phi}^{\prime}}\mathcal{L}_{b})$
9:  end for
10:end for
11:Return: $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ where $\bm{\phi}^{*}\equiv\operatorname{argmin}_{\bm{\phi}}{\mathcal{L}}(\bm{\phi},\bm{\phi}^{\prime})$
In each training iteration (lines 4-10 of Algorithm 2), the model processes a paired sample $(\bm{x}_{-},y_{-})$ and $(\bm{x}_{+},y_{+})$ drawn from the inferior and superior subsets, respectively. The composite loss $\mathcal{L}$, as defined in Eq. 1, is evaluated and backpropagated to update the parameters of both generative models. Through this iterative process, the generative model $\bm{\gamma}$ progressively learns to map inferior solutions to superior ones, guided by the fitness landscape estimated by the surrogate model.
III-D Population Generation
Similar to the reproduction mechanism in traditional EAs, the population generation phase in EvoGO aims to create new candidate solutions based on the current population. However, instead of relying on handcrafted heuristic operators, EvoGO utilizes the learned generative model $\bm{\gamma}$ for solution generation.
Algorithm 3 Population Generation
1:Input: generative model $\bm{\gamma}$ with optimal parameter $\bm{\phi}^{*}$, latest population $\mathcal{D}_{t}$
2:Output: new population $\mathcal{D}_{t+1}$
[…]
4:for each $\left(\bm{x}_{i}^{t},y_{i}^{t}\right)\in\mathcal{D}_{t}$ do
5:  $\bm{x}_{i}^{t+1}\leftarrow\bm{\gamma}(\bm{x}_{i}^{t};\bm{\phi}^{*})$ // “$\bm{\gamma}$” for trained generative model
6:  $y_{i}^{t+1}\leftarrow f(\bm{x}_{i}^{t+1})$ // Function evaluation
7:  $\mathcal{D}_{t+1}\leftarrow\mathcal{D}_{t+1}\cup\left\{\left(\bm{x}_{i}^{t+1},y_{i}^{t+1}\right)\right\}$
8:end for
As illustrated in Algorithm 3, this phase employs the trained generative model $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ to transform inferior solutions into superior ones, thereby generating new candidate solutions from the current population $\mathcal{D}_{t}$. Each generated solution $\bm{x}^{t+1}_{i}$ is evaluated using the true objective function $f$ to obtain its fitness value $y^{t+1}_{i}$, forming the next-generation population $\mathcal{D}_{t+1}$.
Due to the inherently parallelizable nature of the generative model, this step can be fully parallelized, provided that the objective function $f$ supports parallel evaluation—thus significantly improving computational efficiency.
IV Experiments
This section presents a comprehensive evaluation of EvoGO across diverse complex optimization tasks, including low- and high-dimensional numerical benchmarks, classical control problems, and complex high-dimensional robot control tasks. The experiments are designed to assess the effectiveness, efficiency, and scalability of our proposed data-driven evolutionary framework. Specifically, we aim to address the following research questions:
RQ1: How does EvoGO perform in terms of solution quality and convergence speed compared to representative baselines under standard and large population settings?
RQ2: What is the individual contribution of each major component in EvoGO to the overall performance?
RQ3: How does the optimization process of EvoGO evolve over time, and what can be observed from its search behavior through visual analysis?
By systematically exploring these questions, we validate the proposed framework through three groups of experiments: (1) performance and efficiency comparisons under standard and large-scale population regimes; (2) ablation studies to analyze the impact of different components and hyperparameters; and (3) visualizations to reveal the behavioral characteristics of the learned optimization process. These studies collectively demonstrate the scalability, adaptability, and data-efficiency of EvoGO in solving complex optimization problems without reliance on heuristic operations.
IV-A Experimental Setup
Our empirical evaluation covers three categories of optimization challenges: (i) low- and high-dimensional numerical functions, (ii) low- and high-dimensional classical control tasks, and (iii) high-dimensional robotic control tasks. Specifically, the numerical benchmarks include functions from the CEC suite proposed in [90]. The classical control tasks “Pushing” and “Rover” are taken from [91], while “Landing”, “Walker”, and “Ant” are adapted from the Gym library [92]. Detailed descriptions of all benchmark tasks are provided in Supplementary Document C. The maximum number of fitness evaluations (FEs) is set to 1,000 for low-dimensional tasks and 10,000 for high-dimensional tasks.
EvoGO is compared against a comprehensive set of baseline methods, including: (i) Bayesian optimization algorithms with sequential or batch querying, such as Hyperband [93], TuRBO [54], and BOGP (LogEI) [48]; (ii) Tree-structured Parzen Estimator (TPE) [94], a widely used BO method in AutoML; (iii) evolutionary strategies and swarm intelligence algorithms, represented by CMA-ES [27] and PSO [26]; (iv) GPEME [65], a representative surrogate-assisted evolutionary algorithm; (v) Nelder–Mead [95], a classical heuristic method; and (vi) Random Search as a baseline.
All algorithms are evaluated over 20 independent runs with different random seeds on each benchmark. Further implementation details, including optimization settings and training configurations, are provided in Supplementary Document C, while ablation studies on key hyperparameters can be found in Supplementary Document G.
(a) Rosenbrock (5D)
(b) Levy (10D)
(c) Ackley (200D)
(d) Rastrigin (200D)
(e) Landing (12D)
(f) Pushing (14D)
(g) Walker (198D)
(h) Ant (404D)
(i) Plot legend
Figure 3: The convergence curves of EvoGO and the compared baselines in standard settings. The $y$-axes for numerical functions are in logarithmic scale, while those for other tasks are in linear scale since their true optimal is unknown. One-sigma error bars are adopted since two-sigma ones result in excessively large error bars for some algorithms. Please refer to Supplementary Document F for the complete figure.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
(d) Walker (198D)
(e) Ant (404D)
(f) Plot Legend
Figure 4: The convergence curves of EvoGO and the compared baselines (TuRBO, CMA-ES, PSO, and Random Search) in a large-population setting (1000 FEs per generation). All configurations are consistent with those in Fig. 3, except that sequential acquisition algorithms are removed since due to unadjustable population sizes. Please refer to Supplementary Document F for complete figure.
IV-B Convergence Performance across Benchmarks
To address our first research question (RQ1), we evaluate the performance of EvoGO against representative baselines. This set of experiments is designed to assess the final solution quality across a wide range of tasks. By testing on 21 benchmarks spanning numerical functions, classical control, and robot control, with dimensions from 5D to 404D, we aim to demonstrate the method’s broad applicability and effectiveness under both standard and large-population settings.
First, a subset of representative results under standard settings is shown in Fig. 3, while the full results are provided in Supplementary Document F. Under standard population settings, EvoGO consistently outperforms the baselines across most domains, including both low- and high-dimensional numerical functions (Figs. 3(a) to 3(d)), classical control tasks (Figs. 3(e) to 3(f)), and high-dimensional robot control problems (Figs. 3(g) to 3(h)). In total, EvoGO achieves the best performance on 19 out of 21 tasks, 10 of which are statistically significant, and matches the top-performing method on the remaining 9.
To assess scalability under unified conditions, we evaluate several high-dimensional benchmarks using a fixed population size of 1000 across all population-based baselines. As shown in Fig. 4, EvoGO consistently achieves the best performance, demonstrating its ability to efficiently exploit large population diversity in high-dimensional optimization tasks.
We further evaluate EvoGO on the Hopper environment provided by Brax [96], a high-dimensional robot control benchmark that enables massively parallel simulation through GPU acceleration. As shown in Fig. 5, EvoGO achieves the highest final reward among all compared methods, including TPE, CMA-ES, and PPO [97], with respective population sizes of 1000, 1, 100, and 64. Although PPO is an online reinforcement learning algorithm that continuously updates its policy during interaction with the environment, EvoGO achieves comparable or even superior performance. Moreover, EvoGO demonstrates a significant advantage in optimization efficiency: within a time budget of just 500 seconds, it consistently outperforms all baselines—including those running for over an hour. These results highlight the ability of our method to rapidly produce high-quality solutions in complex, hardware-accelerated continuous control tasks.
Figure 5: The rewards of TPE, CMA-ES, PPO, and EvoGO on the Hopper environment provided by Brax. To highlight the advantage in runtime of large population sizes, the rewards are plotted versus the wall-clock time, while the FEs are indicated by different markers. We limited the number of FEs and runtime to 10,000 and one hour respectively for all algorithms.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
Figure 6: The convergence curves of EvoGO and its upper-bound variant “Real Eval”, which directly utilizes real numerical functions instead of surrogate models.
IV-C Convergence Speed across Benchmarks
Further addressing our first research question (RQ1), this section focuses specifically on the convergence speed and efficiency of EvoGO. To this end, our experiments are designed from two angles: first, we measure the number of generations required for convergence on large-scale tasks, and second, we evaluate the wall-clock efficiency within a runtime-constrained environment.
First, as summarized in Table I, EvoGO converges within 10 generations across all large-scale benchmark tasks. In contrast, baseline methods often require tens to thousands of generations to achieve comparable performance. This property is crucial for scenarios with tight runtime budgets but abundant parallel resources, where large populations can be evaluated concurrently.
TABLE I: The number of generations to complete each benchmark by each method. Note: sequential methods are not listed here since they use exactly the same number of generations as FEs.
Benchmark TuRBO CMA-ES PSO Nelder-Mead EvoGO
Numerical 10$\sim$100 83$\sim$526 25$\sim$50 956$\sim$9395 10
Landing 10 91 25 823$\sim$901 10
Pushing 10 91 25 862$\sim$933 10
Rover 100 625 25 3163$\sim$3216 10
Walker 100 526 50$\sim$51 1833$\sim$1839 10
Ant 100 455 24$\sim$25 1500$\sim$1503 10
Second, the runtime-constrained setting of Brax allows us to evaluate the wall-clock efficiency of each method under the same budget of 10,000 function evaluations and one hour of runtime. As illustrated in Fig. 5, EvoGO rapidly approaches a reward of 900 within the first 100 seconds, while other methods take significantly longer to reach similar performance—or fail to reach it at all. In particular, EvoGO achieves up to 134×134$\times$134 × speedup over CMA-ES and 50×50$\times$50 × over TPE in terms of reward per second. This demonstrates the strong convergence speed of our framework when leveraging large population sizes and parallel evaluations.
IV-D Ablation Studies
To answer our second research question (RQ2), we conduct a comprehensive ablation study to dissect EvoGO and quantify the contribution of its key components. These experiments are designed to isolate the impact of elements like the surrogate model, infill criterion, and generative architecture on overall performance.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
(d) Walker (198D)
(e) Ant (404D)
(f) Plot Legend
Figure 7: The convergence curves of EvoGO and its ablations. “EvoGO” represents the original method; “Single Net” excludes the generative model $\bm{\gamma}^{\prime}$ and related loss terms; “Cycle GAN” adapts the Cycle GAN; “EvoGO-MLP” utilizes MLP for surrogate modeling; “EvoGO-LCB” employs the LCB infill criteria. Please refer to Supplementary Document G for the complete figure.
1) Upper Bound of the Generative Model
We begin by investigating the maximum potential of our framework to understand its upper performance bound. This experiment assesses the effectiveness of the overall architecture, independent of any surrogate model error. To achieve this, we replace the surrogate model’s predictions with ground-truth function evaluations in the optimization loss (“Real Eval”). This design serves two purposes: first, it reveals the high potential of our composite generative structure and optimization strategy, and second, it suggests that performance could be further enhanced with more accurate surrogate models. As presented in Fig. 6, the “Real Eval” method shows robust and consistent performance across various numerical functions, illustrating the effectiveness of EvoGO’s overall architecture. The great potential of the proposed composite generative model structure and the optimization loss is well demonstrated by the results, which can be considered as an upper bound of EvoGO.
2) Alternative Surrogate Models
We explore the framework’s modularity and sensitivity to the choice of surrogate model. This experiment is designed to understand how performance is influenced by different modeling assumptions. We replace the default Gaussian Process (GP) with a Multi-Layer Perceptron (MLP) [98] (“EvoGO-MLP”), a common alternative that lacks inherent uncertainty quantification. In both variants, the standard deviation and covariance functions are removed from the acquisition loss (Eq. 11). This allows us to assess the importance of the GP’s probabilistic nature and to demonstrate the framework’s adaptability. As shown in Fig. 7, the “EvoGO-MLP” struggled with some numerical functions, likely due to overfitting and lacking the variance estimation provided by GPs. However, it exhibited impressive results on complex scenarios like the Ant benchmark. These findings highlight EvoGO’s adaptability and its potential utility with different surrogate modeling methods.
We examine the modular structure of the framework and its performance under different surrogate modeling assumptions. We replace the default Gaussian Process (GP) with a Multi-Layer Perceptron (MLP) [98] (referred to as “EvoGO-MLP”), a common alternative that lacks inherent uncertainty quantification. In both variants, the standard deviation and covariance components are removed from the acquisition loss (Eq. 11), enabling an evaluation of the role of probabilistic modeling in GP-based surrogates and demonstrating the framework’s flexibility. As shown in Fig. 7, “EvoGO-MLP” underperforms on certain numerical benchmarks, likely due to overfitting and the absence of variance estimation available in GPs. However, it performs well on complex control tasks such as the Ant benchmark. These results demonstrate that EvoGO remains effective across different surrogate modeling choices while highlighting the value of uncertainty modeling in challenging scenarios.
3) Alternative Infill Criteria
We replace our tailored loss (Eq. 11) with one derived from the widely used Lower/Upper Confidence Bound (LCB/UCB) infill criteria (“EvoGO-LCB”) to validate the effectiveness of our custom optimization loss. By comparing the performance of variant against the original EvoGO, we can determine the specific benefits of criterion and test the framework’s flexibility with alternative infill criteria. Specifically, the optimization loss of the LCB/UCB infill criteria can be written as:
$\mathcal{L}_{\text{opt}}^{\text{LCB}}=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2},$ (12)
where the derivation is detailed in Supplementary Document A. As shown in Fig. 7, the “EvoGO-LCB” variant generally performed worse than the original EvoGO. However, it still achieved competitive results on the robot control and some numerical benchmarks. This not only suggests that the original optimization loss of EvoGO (Eq. 11) is more robust across a wider range of tasks but also highlights the adaptability of the proposed EvoGO.
Generation 1
Generation 3
Generation 5
PH
PH
Fitness
Figure 8: Visual representation of vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$ on the Ackley function landscape. The vectors were generated by the trained generative model across different generations. They illustrate the transformation from input solutions $\bm{x}_{\text{in}}$ to the corresponding output solutions $\bm{x}_{\text{out}}$. The central star indicates the global optimum, while the dashed rectangles highlight regions covered by the solutions generated at different generations. For clarity, any landscape shifting and rotation have been omitted from this visualization.
4) Without Surrogate Model
In this section, we evaluate whether the surrogate-guided acquisition process is essential, or if the generative component alone can effectively guide the search. To this end, we test two variants: one that completely removes the surrogate model and its acquisition loss (“No Surrogate”), and another that replaces it with a standard adversarial loss from a Cycle GAN framework [75] (“Cycle GAN”). Specifically, the latter incorporates the data setup described in Algorithm 1, and replaces GP posterior and related acquisition loss (Eq. 11) with discriminative model and adversarial loss, respectively. Moreover, to train the “Cycle GAN” variant, we adopted the gradient inversion technique, which is not present in EvoGO. This allows us to evaluate the necessity of the surrogate model for balancing exploration and exploitation. As shown in Fig. 7, these models generally showed diminished effectiveness, likely due to increased susceptibility to mode collapse. In contrast, EvoGO adaptively balances exploration and exploitation, thus avoiding such pitfalls.
5) Unpaired Generative Modeling
Finally, we investigate the importance of the paired generative architecture, a core structural choice in EvoGO. To do this, we test a “Single Net” variant with only one generative model, removing the losses that enforce consistency between the pair (Eq. 5). Specifically, it eliminates the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5) and half of the similarity loss $\mathcal{L}_{\text{sim}}$ (Eq. 6) in Eq. 7. Furthermore, the standard deviation loss $\mathcal{L}_{\text{std}}$ in Eq. 11 is also removed in the “Single Net” variant. This experiment is crucial for validating our hypothesis that the paired design is key to preventing mode collapse and ensuring robust performance. The results shown in Fig. 7 confirms that this variant underperformed significantly across most benchmarks, underscoring the critical role of paired modeling in maintaining population diversity and enhancing overall system performance.
6) Extended Ablations
Additional ablation studies are detailed in Supplementary Document G. These studies examined variations in the neural network architectures and hyperparameter configurations, further elucidating the robustness and adaptability of EvoGO across various settings. Specifically, we explored the impact of different generative model layers and the width of them, noticing that the generative model’s architecture has little effect on EvoGO’s performance, as shown in Fig. S4. Moreover, we also investigated the impact of the reconstruction loss scale $\lambda_{1}$, the data split factor $\eta$, and the sliding window factor $\epsilon$, all of which were found to have small to little impact on the performance of EvoGO, as shown in Figs. S5 to S7.
IV-E Search Dynamics
Finally, to answer our third research question (RQ3) and gain deeper insights into the search dynamics of EvoGO, we conducted a visualization study on a toy problem. Specifically, Fig. 8 investigates the search dynamics of EvoGO by applying it to a two-dimensional Ackley function with a population size of 100. Here, we tracked the transitions from input solutions $\bm{x}_{\text{in}}$ to generated output solutions $\bm{x}_{\text{out}}=\bm{\gamma}(\bm{x}_{\text{in}})$, captured across various generations from the trained generative model $\bm{\gamma}$. Arrows in the visualization represent vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$, and the colors of these arrows reflect their norm-2 lengths, as described in the corresponding plot legends. This visual representation clearly demonstrates EvoGO’s ability to precisely estimate both the direction $(\bm{x}_{\text{out}}-\bm{x}_{\text{in}})/\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ and the magnitude $\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ of steps towards potential optima. This capability not only enhances optimization efficiency but also significantly mitigates mode collapse. Furthermore, $\bm{\gamma}$’s learning progress is illustrated through the grayscale contours of the optimization landscapes captured by the solutions at different generations. Throughout the search process, $\bm{\gamma}$ adaptively hones in on specific regions, showing a refined focus that enhances its search efficacy. This adaptiveness suggests EvoGO’s potential for achieving superlinear convergence given ample data, which is accentuated by the dashed rectangles that delineate the boundaries of solutions generated at different generations.
V Conclusion
This paper presents Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework designed to address three critical limitations in existing generative model-based evolutionary algorithms: reliance on handcrafted heuristics, objective misalignment between generative models and evolutionary goals, and underutilization of available data. By eliminating traditional search operators and adopting globally consistent training objectives, EvoGO enables end-to-end learning for optimization with enhanced generality, alignment, and data efficiency. The proposed framework introduces a paired generative architecture that integrates forward and inverse models to produce high-quality candidate solutions, effectively replacing conventional crossover and mutation. A pairwise data construction strategy is employed to enhance training diversity under limited evaluation budgets. Meanwhile, a composite loss function that integrates reconstruction, distributional alignment, and directional guidance ensures efficient model convergence toward promising regions. Extensive experiments across numerical, control, and robotic tasks demonstrate the superior performance of EvoGO over state-of-the-art baselines. It achieves convergence within 10 generations and delivers up to a 134× speedup on complex tasks through GPU-parallelized inference. These findings highlight the potential of fully learnable evolutionary frameworks to generalize across domains without manual design. Future work will focus on enhancing performance under extremely low-data regimes and providing theoretical insights into convergence behavior and optimization dynamics.