### v2

# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

Abstract
Recent advances in data-driven evolutionary algorithms (EAs) have demonstrated the potential of leveraging historical data to improve optimization accuracy and adaptability. Despite these advancements, existing methods remain reliant on handcrafted process-level operators. In contrast, Evolutionary Generative Optimization (EvoGO) is a fully data-driven framework designed from the objective level, enabling autonomous learning of the entire search process. EvoGO streamlines the evolutionary optimization process into three stages: data preparation, model training, and population generation. The data preparation stage constructs a pairwise dataset to enrich training diversity without incurring additional evaluation costs. During model training, a tailored generative model learns to transform inferior solutions into superior ones. In the population generation stage, EvoGO replaces traditional reproduction operators with a scalable and parallelizable generative mechanism. Extensive experiments on numerical benchmarks, classical control problems, and high-dimensional robotic tasks demonstrate that EvoGO consistently converges within merely 10 generations and substantially outperforms a wide spectrum of optimization approaches, including traditional EAs, Bayesian optimization, and reinforcement learning based methods. Code is available at: https://github.com/EMI-Group/evogo
I Introduction
Many real-world optimization tasks involve objectives that can be accessed only through simulations, legacy software, or physical experiments [1, 2]. Such problems are typically non-analytical, noisy, and gradient-free, making them challenging for conventional optimization techniques. In these scenarios, Evolutionary Algorithms (EAs) provide a robust and widely applicable framework due to their global search capability, insensitivity to differentiability, and ease of parallelization. They have achieved success in diverse fields such as hyperparameter tuning, biomedical engineering, and aerospace design [3]. Nevertheless, while EAs offer a practical starting point for complex black-box problems, their classical forms remain limited in capturing structure or reusing information across generations [4, 5, 6, 7].
To enhance efficiency and stability, subsequent research reformulated EAs from an operator-based to a distribution-based paradigm. Representative examples include Evolution Strategies (ES) [8] and Estimation of Distribution Algorithms (EDAs) [9], which replace handcrafted variation operators with statistical modeling and parameterized sampling. This allows population updates through distributional shifts rather than discrete manipulations, while maintaining desirable invariance properties such as rotation or scale invariance [10]. By estimating gradients in the distribution space, these methods introduce mechanisms like natural gradient ascent, covariance matrix adaptation [11], and factorized or low-rank modeling to better capture dependencies among decision variables[12]. Such strategies effectively adjust the sampling distribution using the data collected from previous evaluations, thereby improving sample efficiency and convergence stability [13]. However, their optimization still focuses on fitting the sampling distribution itself, rather than learning a coherent decision process that couples evaluation and generation in a data-driven manner.
To further enhance the consistency, adaptivity, and data efficiency of evolutionary optimization, subsequent studies have shifted attention from merely updating the sampling distribution to learning the evaluation process itself [14]. This line of research gave rise to data-driven evolutionary frameworks, particularly surrogate-assisted ones, where the objective landscape is approximated through learned evaluators or ranking models trained on historical samples [15, 16]. By replacing direct evaluations with predictive or ordinal surrogates, these methods achieve more consistent and controllable optimization dynamics while significantly reducing computational cost [17, 18, 19]. Fundamentally, the evaluation step becomes a learned component that embeds accumulated experience into the search process, thereby narrowing the gap between evolutionary optimization and data-driven inference. However, their reliance on imperfect surrogates still introduces model bias, and repeated exploitation of surrogate predictions can lead to premature convergence and loss of exploration diversity.
To mitigate these limitations, generative modeling-based evolutionary algorithms (GMEAs) [20, 21, 22] have emerged as a promising extension of data-driven evolutionary computation. Instead of learning to approximate the evaluation process, GMEAs focus on learning the generation process itself, representing the search dynamics through parameterized probabilistic models. By coupling the exploratory nature of evolutionary search with the expressive power of generative modeling, GMEAs establish a positive feedback loop between solution generation and model learning  [23, 24, 25, 26]. GMEAs are particularly suited for complex black-box optimization problems that exhibit latent dependencies, implicit constraints, or simulation-based evaluations, where gradients or analytical forms are unavailable. Through data-driven modeling of the search distribution, GMEAs achieve a more balanced trade-off between exploration and exploitation, effectively alleviating the model bias and premature convergence issues often observed in surrogate-assisted approaches.
However, existing GMEAs are still at an early stage and face several fundamental challenges. First, their integration with evolutionary workflows often relies on manually designed operators and procedural rules, thereby embedding task-specific heuristics into the algorithm rather than achieving automatic coordination between the generative model and the evolutionary process. Second, the training objectives of current generative models are typically inherited from domains such as image or text generation, leading to misalignment with optimization goals. Third, most methods depend on limited population data, which hampers training stability and generalization under tight evaluation budgets. These limitations highlight the need for a fully data-driven evolutionary framework that replaces handcrafted heuristics with end-to-end, learning-based mechanisms. To this end, we propose Evolutionary Generative Optimization (EvoGO), a unified framework that achieves evolutionary optimization entirely through generative learning. Our key contributions are summarized as follows:
  • We design a paired generative architecture in which forward and inverse models collaborate to produce high-quality candidate solutions. This setup enables direct mapping from historical populations to new individuals, entirely replacing traditional crossover and mutation heuristics.
  • We introduce a tailored loss function that combines reconstruction accuracy, distributional alignment, and directional guidance. This composite objective explicitly steers the generative process toward promising regions of the search space.
  • We develop a learning-based sample synthesis strategy that generates informative training data from limited fitness evaluations. This mechanism supports stable model training even under small population sizes and tight evaluation budgets.
The remainder of the paper is organized as follows. Section II reviews existing approaches for complex optimization and discusses their relationships to EvoGO. Section III introduces the proposed model architecture, loss function, and learning strategies. Section IV presents a comprehensive empirical evaluation, including performance comparisons and ablation studies. Finally, Section V concludes the paper.
II Related Work
II-A Non-Data-Driven Approaches
EAs represent the most prevalent non-data-driven paradigm for complex optimization. These approaches rely on fixed heuristics and do not leverage learned, data-driven models to guide search. EAs can be broadly categorized based on their mechanism: (i) methods that generate solutions via handcrafted search operators, and (ii) methods that probabilistically model the distribution of promising solutions to guide sampling.
Methods based on handcrafted operators, such as genetic algorithms [27], genetic programming [28], differential evolution [29], and particle swarm optimization [30], constitute the first category. These algorithms rely on fixed, manual operators (e.g., crossover or mutation) to explore the search space. Despite their widespread adoption, their performance is heavily dependent on domain knowledge and requires substantial manual tuning, which restricts both flexibility and scalability. To address these limitations, the second category attempts to guide the evolutionary process by learning distributions over promising solutions.
The second category comprises Distribution-based EAs, which aim to improve search efficiency by modeling the distribution of successful solutions. This category includes Evolution Strategies (ES) and Estimation of Distribution Algorithms (EDAs). ES, originating from early stochastic optimization studies [31], achieved a major advancement with Covariance Matrix Adaptation ES (CMA-ES), which adaptively adjusts the sampling distribution based on population statistics [11] and was later extended to a fully derandomized formulation [32]. More recently, ES variants such as OpenAI-ES [33] have demonstrated strong scalability and parallelism in large-scale machine learning tasks. In parallel, EDAs explicitly incorporate statistical modeling into evolutionary search by learning the distribution of promising solutions. Early examples include Population-Based Incremental Learning (PBIL) [13] and the Univariate Marginal Distribution Algorithm (UMDA) [12]. Subsequent methods, such as the extended compact genetic algorithm [34] and the linkage tree genetic algorithm [35], capture variable dependencies through more expressive probabilistic models.
Despite their conceptual elegance, both ES and EDA methods remain fundamentally constrained by the non-data-driven paradigm. The underlying update and sampling procedures are still manually crafted and adhere to rigid algorithmic rules, which severely restricts their adaptability across diverse or complex optimization scenarios [8, 9]. To address this lack of flexibility, some research efforts have explored incorporating reinforcement learning (RL) into EAs to automate the selection or configuration of evolutionary operators [36, 37, 38, 39, 40]. These methods typically employ external RL agents to dynamically choose strategies based on past performance. However, while this integration offers a degree of adaptivity, the learning process is still external to the core optimization loop, often incurring substantial computational overhead and poor sample efficiency [41]. Consequently, these approaches do not fundamentally depart from the heuristic nature of traditional EAs, and the search dynamics remain largely heuristic in nature.
II-B Data-Driven Apporoaches
Traditional EAs primarily rely on handcrafted heuristics and static update rules. However, recent research has increasingly shifted toward data-driven paradigms that exploit information accumulated during the optimization process. These methods employ machine learning techniques to enhance the efficiency, scalability, and adaptability of evolutionary optimization. Consequently, these approaches are generally categorized into three representative categories: Bayesian Optimization (BO), Surrogate-Assisted Evolutionary Algorithms (SAEAs), and Generative Modeling-based Evolutionary Algorithms (GMEAs).
1) Bayesian Optimization
Bayesian Optimization (BO) traditionally employs Gaussian processes to model the objective function and constructs acquisition functions to determine promising query points [42]. Substantial efforts have been made to enhance the scalability of BO to address performance challenges in large-scale and high-dimensional settings [43, 44, 45, 46, 47, 48, 49, 50]. Despite these advances, significant computational overhead persists, especially when BO is applied to complex tasks on modern high-throughput computing platforms.
To address this persistent computational overhead, researchers have proposed multi-point acquisition strategies designed to enable efficient batch querying [51, 52, 53, 54]. Among these approaches, Thompson Sampling (TS) [55] has gained particular popularity for use in parallel BO schemes [56, 57, 58, 59]. For instance, Eriksson et al. integrated TS with a trust region method to achieve scalable BO for constrained and unconstrained problems, supporting batch sizes up to 100. Nevertheless, in the majority of cases, these batch sizes remain relatively modest, thereby resulting in the underutilization of modern computational resources.
In parallel with multi-point methods, Artificial Neural Networks (ANNs) have also been explored to improve the batch efficiency of BO [60, 61, 62]. These approaches, however, often repurpose the surrogate model as a binary classifier rather than directly generating new candidate solutions. For instance, density-ratio-based methods were proposed by Tiao et al. and Oliveira et al., while Swersky et al. adopted reinforcement learning to train a selection policy for discrete domains. Despite their contributions, these methods still fundamentally rely on direct pointwise optimization for generating new queries [63], a mechanism that ultimately hinders the full scalability of BO.
2) Surrogate-Assisted Evolutionary Algorithms
Surrogate-Assisted Evolutionary Algorithms (SAEAs) aim to enhance sample efficiency in optimization problems where fitness evaluations involve costly simulations [64, 7]. Early frameworks in model-based evolutionary computation introduced approximation models to reduce evaluation cost [65, 66]. Traditional SAEAs typically employ simple surrogate models, such as Radial Basis Functions (RBFs) and Support Vector Machines (SVMs) [67]; however, these models are limited in capturing the complexity of high-dimensional and nonlinear search landscapes. The subsequent introduction of Gaussian Processes (GPs) alleviated some of these constraints by incorporating uncertainty estimation into the infill process. For instance, Liu et al. [68] combined GPs with standard EA operators to effectively balance exploration and exploitation.
However, since only a limited number of training samples are available during the optimization process, a single surrogate model often struggles to generalize across heterogeneous problem domains or adapt to different search stages. To address this issue, recent studies have proposed multi-surrogate frameworks that better capture diverse landscape features and mitigate model bias. These frameworks can be broadly classified into two categories: hierarchical-model SAEAs and ensemble-model SAEAs. The hierarchical approach combines global and local models to improve scalability [16, 69, 70], whereas ensemble paradigms integrate multiple predictors to enhance accuracy and reduce overfitting [71, 72]. In parallel, adaptive-model SAEAs have been developed to automatically select appropriate surrogate models or infill criteria according to evolving problem characteristics [18, 19, 73]. Despite these advances, most SAEAs still rely heavily on surrogate predictions for candidate selection, which can be computationally demanding and prone to model bias [73].
3) Generative Modeling-based Evolutionary Algorithms
Generative Modeling-based Evolutionary Algorithms (GMEAs) aim to learn and adapt the sampling distribution from data to guide the evolutionary search process. Current research in this area mainly follows two directions: (i) leveraging unsupervised learning for variable grouping and dimensionality reduction [74, 75, 20], and (ii) employing supervised or reinforcement learning to guide the directional improvement of candidate solutions [76, 77, 78, 21, 22].
The first line of work trains generative models to discover latent structures within the search space, thereby facilitating more efficient variation. For example, Pan et al. [20] proposed a manifold-learning-based approach that restricts mating operations to low-dimensional subspaces, improving optimization efficiency.
The second line focuses on guiding evolutionary search using data-driven feedback. He et al. [78] introduced a GAN-based multi-objective EA in which a discriminator distinguishes real from synthetic populations to train a generator that produces new candidates. Similarly, Liu et al. [21] employed reinforcement learning to adaptively select variation operators based on individual context and historical performance. While GMEAs show strong potential, several open challenges remain. These include their continued reliance on manually designed heuristic components, the difficulty of seamlessly integrating generative models into the evolutionary loop, and ambiguities in defining effective training objectives.
II-C Discussion
The evolution from non-data-driven to data-driven paradigms in evolutionary computation reflects a broader shift toward leveraging information accumulated during optimization. Despite significant progress across BO, SAEA, and GMEAs, each paradigm exhibits distinct advantages and limitations that restrict its general applicability.
BO offers strong theoretical foundations and high sample efficiency but suffers from poor scalability due to the cubic complexity of Gaussian process inference and reliance on pointwise query generation. In contrast, SAEA improves efficiency by introducing learned evaluators to approximate the fitness landscape; however, its focus on local modeling often results in bias and premature convergence. Finally, GMEA emphasizes learning the distribution of promising solutions from data, enabling better structural modeling and diversity maintenance. Nevertheless, current GMEA implementations still rely on manually designed heuristic components and lack effective coupling between the generation and evaluation phases.
To address the existing limitations identified in current approaches, the research imperative is to develop a unified, end-to-end learning framework that seamlessly integrates representation learning, solution generation, and fitness evaluation within a cohesive pipeline. Such a framework must satisfy key design requirements, specifically seeking to (i) eliminate reliance on ad hoc or handcrafted coordination mechanisms, (ii) ensure rigorous alignment between the learning objectives of each internal module and the global optimization goal, and (iii) effectively leverage all available information to maximize data efficiency and adaptability.
III Method
This section introduces the overall framework of EvoGO, followed by detailed descriptions of each key component.
Figure 1: The overall framework of EvoGO. The historical dataset $\mathcal{D}$ is composed of $\{\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}\}$, where each $\mathcal{D}_{t}=\left\{(\bm{x}_{t,i},y_{t,i})\mid 1\leq i\leq N\right\}$ , with $\bm{x}_{t,i}$ and $y_{t,i}$ denoting the population and its fitness, respectively. In data preparation phase, the historical population is first selected to form the training set of current generation and augmented by an optional learning-based data augmentation strategy. Then, it is separated into the superior and inferior groups with smaller and larger fitnesses, respectively, before a Cartesian product is applied to generate the paired training dataset. In model training phase, the framework first trains a surrogate model solely for landscape estimation. Then, the surrogate model is frozen and combined with a pair of generative models to form a composite model for effective generative model training. Later, the composite model is trained via the proposed training loss tailored for comoplex optimization. In population generation phase, the inferior-to-superior generative model is used to generate the new solutions in parallel.
III-A Overall Framework
Algorithm 1 Data Preparation
1:Input: historical data $\mathcal{D}$, dataset split factor $\eta\in(0,1)$, sliding window factor $\epsilon>=0$, data augmentation threshold NN, and factor $\rho$
2:Output: $\mathcal{D}_{\text{SM}}$ for surrogate model training, $\mathcal{D}_{\text{GM}}$ for generative model training
3: /************** Elite Selection ***************/
[…]
6: /******** Data Augmentation (Optional) ********/
7:if $|\mathcal{D}_{0}|<N$ then
8:  $\tilde{f}\leftarrow$ trained surrogate model on $\mathcal{D}_{\text{SM}}$
9:  $\bm{\mathcal{X}}^{\prime}\leftarrow$ randomly generate $\rho\cdot|\mathcal{D}^{\prime}|$ solutions according to the distribution of solutions in $\mathcal{D}^{\prime}$
10:  $\mathcal{D}^{\prime}\leftarrow\{(\bm{x},\mu_{t}(\bm{x}))|\bm{x}\in\bm{\mathcal{X}}\}$ where $\mu_{t}(\bm{x})=\mathbb{E}[\tilde{f}(\bm{x})]$
11:end if
12: /*********** Paired Set Construction ***********/
[…]
As illustrated in Fig. 1, EvoGO comprises three key, sequentially executed phases: data preparation, model training, and population generation. To streamline the evolutionary process and eliminate reliance on complex handcrafted rules, EvoGO adopts an end-to-end generative paradigm, replacing the heuristic components commonly used in traditional GMEAs.
This generative architecture directly addresses two core challenges inherent in existing GMEAs: data scarcity and objective misalignment. Specifically, the data preparation phase incorporates a data augmentation strategy to mitigate the risk of overfitting caused by limited population data, thereby enhancing training diversity without introducing additional evaluation costs. Furthermore, in the model training phase, a composite model architecture and a tailored loss function are employed to align the generative objectives with the goals of evolutionary optimization, thereby resolving the objective misalignment issue. The time and space complexity of EvoGO are analyzed in detail in Supplementary Document A.
III-B Data Preparation
The data preparation phase is pivotal for constructing high-quality and well-structured training data that is sufficient for both the surrogate model and the generative model, particularly under limited evaluation budgets. Achieving high generalization capabilities for the generative model requires rich, diverse input data, whereas the surrogate model demands highly promising samples for accurate fitness estimation. Consequently, this phase is designed to manage the trade-off between diversity and fidelity to maximize data utility.
As detailed in Algorithm 1, the data preparation process comprises the following three steps:
  • Elite Selection: A high-quality dataset $\mathcal{D}_{\text{SM}}$ is constructed by selecting individuals with the highest fitness values from all historical populations $\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}$. This process ensures accurate surrogate modeling by focusing on representative and promising solutions.
  • Data Augmentation (Optional): When the evaluation budget is limited, requiring the algorithm to operate with a small population, a learning-based data augmentation strategy is adopted. Specifically, a Variational AutoEncoder (VAE) is used to generate diverse synthetic samples that resemble the current population. This process enriches the generative model dataset $\mathcal{D}_{\text{GM}}$, thereby alleviating the data scarcity issue under restricted evaluations.
  • Paired Set Construction: Instead of enlarging the dataset further, this step enhances data utilization by constructing informative sample pairs. The surrogate model estimates the fitness of all candidates, after which the samples are ranked and divided into superior ($\mathcal{D}_{\text{GM}^{+}}$) and inferior ($\mathcal{D}_{\text{GM}^{-}}$) subsets according to a predefined ratio $\eta$. The Cartesian product of the two subsets forms the paired training set $\mathcal{D}_{\text{GM}}$, enabling the generative model to learn both effective descent directions and adaptive step magnitudes through relational supervision.
III-C Model Training
The model training phase focuses on establishing a composite model architecture integrated with a dedicated composite loss function. This architecture enables directional learning and optimization-aware training through an online process, where lightweight training is conducted during each evolutionary generation. Although the composite structure involves three model components, only the forward generative model $\bm{\gamma}$ is ultimately used during the population generation stage. The subsequent text details the overall training procedure, beginning with the model architecture and followed by the formal loss formulation.
1) Model Architecture
Inspired by Generative Adversarial Networks (GANs) [79] and CycleGANs [80], we develop a composite-model-based GMEA framework (see Fig. 2) that integrates a surrogate model $\tilde{f}$ with two complementary generative components. These components are a forward generative model $\bm{\gamma}$, which maps inferior solutions to superior ones, and a pseudo-inverse model $\bm{\gamma}^{\prime}$, which performs the reverse mapping. Unlike conventional GAN-like architectures that rely on discriminators, the proposed framework employs the surrogate model to enforce optimization-aware learning, thereby aligning generative behaviors with the underlying objective landscape.
Given an evolving population, the composite architecture is designed to generate new solutions that possess improved function values. The dual generative models, $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$, implement the generative loss $\mathcal{L}_{\text{gen}}$ (Eq. 2 and Eq. 7), whereas the surrogate model $\tilde{f}$ addresses the optimization loss $\mathcal{L}_{\text{opt}}$. Specifically, $\bm{\gamma}$ processes inferior solutions to generate estimated superior ones, and $\bm{\gamma}^{\prime}$ operates in reverse, both facilitating the computation of the similarity loss $\mathcal{L}_{\text{sim}}$.
The integrated architecture further facilitates the transfer of optimization landscape estimation from the surrogate model to the dynamic generative components. Initially, the surrogate model $\tilde{f}$ is established and remains constant throughout the training process. Subsequently, the solutions generated by $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are reconstructed by cycling back to each other, which facilitates the computation of the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5). Crucially, the generated solutions ($\bm{\gamma}(\bm{p})$ and $\bm{\gamma}^{\prime}(\bm{q})$) are also evaluated against the surrogate, contributing to the optimization loss $\mathcal{L}_{\text{opt}}$. This configuration allows for the transfer of optimization landscape information via back-propagation, thereby enabling $\bm{\gamma}$ to autonomously generate optimized populations.
In summary, the architecture synergistically integrates a static surrogate model $\tilde{f}$ with a dynamic generative model $\bm{\gamma}$ and its pseudo-inverse $\bm{\gamma}^{\prime}$. $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are specifically trained for generating and reconstructing the superior and inferior solutions, respectively. Conversely, $\tilde{f}$ focuses exclusively on the estimation of the optimization landscape. This synergistic training substantially enhances the practical ability of $\bm{\gamma}$ to generate optimized solutions.
2) Loss Function
As illustrated in Fig. 2, we define a composite loss function that integrates two principal components: the generative loss $\mathcal{L}_{\text{gen}}$ and the optimization loss $\mathcal{L}_{\text{opt}}$:
$\mathcal{L}=\mathcal{L}_{\text{gen}}+\lambda\mathcal{L}_{\text{opt}},$ (1)
where $\lambda$ is a trade-off parameter that balances generative learning and optimization guidance.
The generative loss $\mathcal{L}_{\text{gen}}$ comprises two core components. The first is the similarity loss $\mathcal{L}_{\text{sim}}$, which encourages consistent latent representations between paired samples. The second is the reconstruction loss $\mathcal{L}_{\text{rec}}$, which enforces structural consistency in the generated outputs:
$\mathcal{L}_{\text{gen}}=\mathcal{L}_{\text{sim}}+\lambda_{1}\mathcal{L}_{\text{rec}}.$ (2)
The overall optimization loss $\mathcal{L}_{\text{opt}}$ is formulated as the combination of two components: the expectation loss $\mathcal{L}_{\text{mean}}$, which penalizes poor average fitness among generated samples, and the variance loss $\mathcal{L}_{\text{std}}$, which promotes diversity:
$\mathcal{L}_{\text{opt}}=\mathcal{L}_{\text{mean}}+\lambda_{2}\mathcal{L}_{\text{std}},$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the contributions of the reconstruction and variance terms, respectively.
The similarity loss $\mathcal{L}_{\text{sim}}$ is conventionally employed in generative learning paradigms to ensure that generated outputs closely resemble the target samples [81, 80, 82, 83]. Specifically, for the generative model $\bm{\gamma}$, which transforms a source solution $\bm{p}$ to a target solution $\bm{q}$, we adopt a mean squared loss:
$\mathcal{L}_{\text{sim}}=\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}.$ (4)
This aligns with conventional generative paradigms, where models learn mappings from latent or random variables to target domains [79, 84, 85]. However, when applied to optimization problems, such models often struggle with the intrinsic trade-off between exploration and exploitation, especially within the EA framework [5, 86, 87]. Alternatively, domain adaptation models, which transform solutions between source and target domains [88, 89, 80, 90], have shown promising results in certain optimization scenarios [87, 91]. Nonetheless, without reconstruction constraints, these models are susceptible to mode collapse [81, 82, 83], thereby compromising training stability. To mitigate this issue, we introduce a paired generative model $\bm{\gamma}^{\prime}$ that serves as a pseudo-inverse of $\bm{\gamma}$. This yields a reconstruction loss:
$\mathcal{L}_{\text{rec}}=\|\bm{\gamma}^{\prime}(\bm{\gamma}(\bm{p}))-\bm{p}\|^{2}+\|\bm{\gamma}(\bm{\gamma}^{\prime}(\bm{q}))-\bm{q}\|^{2},$ (5)
thereby leading to the similarity loss symmetrized as:
$\mathcal{L}_{\text{sim}}=\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}+\|\bm{\gamma}^{\prime}(\bm{q})-\bm{p}\|^{2}.$ (6)
The generative loss $\mathcal{L}_{\text{gen}}$, particularly the reconstruction term, plays an essential role in stabilizing training and preserving semantic fidelity. This mechanism is crucial in optimization contexts, where the model must accurately learn transformations from inferior to superior regions of the search space without drifting off-manifold. To ensure this, the weight $\lambda_{1}$ associated with the reconstruction loss is set to dominate the loss function, emphasizing faithful reconstruction over mere proximity.
By integrating the similarity and reconstruction losses for the bidirectional models $\{\bm{\gamma},\bm{\gamma}^{\prime}\}$, we obtain the full generative loss:
$\mathcal{L}_{\text{gen}}$ $=\left(\|\bm{\gamma}(\bm{p})-\bm{q}\|^{2}+\|\bm{\gamma}^{\prime}(\bm{q})-\bm{p}\|^{2}\right)$
$\quad+\lambda_{1}\left(\|\bm{\gamma}^{\prime}(\bm{\gamma}(\bm{p}))-\bm{p}\|^{2}+\|\bm{\gamma}(\bm{\gamma}^{\prime}(\bm{q}))-\bm{q}\|^{2}\right).$ (7)
This formulation adheres to the principles of domain-adaptive generative learning, where $\bm{p}$ and $\bm{q}$ are samples from source (inferior) and target (superior) domains, respectively. The dataset construction process is detailed in Section III-B, and an alternative interpretation of Eq. 7 from a probabilistic perspective is provided in Supplementary Document B.
Integrating the generative model into the EA framework necessitates transforming traditional optimization objectives into suitable loss function terms for a complete data-driven process. This requirement is analogous to the infill criteria used in SAEAs, where the surrogate model selects the most promising solutions from a population of candidates [73]. Such selection strategies have been widely studied in the literature. One of the most common approaches is to use the Upper/Lower Confidence Bound (UCB/LCB) [43] as the infill criteria [68]:
$\alpha(\bm{p})=\mathbb{E}\left[\tilde{f}(\bm{p})\middle|\bm{p}\right]+\lambda_{2}\mathbb{V}\left[\tilde{f}(\bm{p})\middle|\bm{p}\right]^{1/2},$ (8)
Traditional Surrogate-Assisted Evolutionary Algorithms (SAEAs) utilize the infill criteria, $\alpha(\bm{p})$, to guide the selection of candidate solutions $\bm{p}$ for high-fidelity evaluation. Specifically, $\tilde{f}$ denotes the surrogate of the objective function being minimized, and $\lambda_{2}$ is a hyperparameter defined in Eq. 3. All candidate solutions are evaluated using the criterion established in Eq. 8 to identify those with minimal $\alpha(\bm{p})$ values, which are subsequently passed to the true objective function for final evaluation.
Figure 2: Overview of the model architecture and composite loss functions $\mathcal{L}$ in EvoGO. The top, middle, and bottom rows illustrate the computational workflows for the similarity loss $\mathcal{L}_{\text{sim}}$, the reconstruction loss $\mathcal{L}_{\text{rec}}$, and the optimization loss $\mathcal{L}_{\text{opt}}$, respectively. The input pair $(\bm{p},\bm{q})$ corresponds to solutions with inferior and superior fitness, respectively.
However, the conventional application of infill criteria, such as UCB or LCB, is not directly applicable to our proposed generative model $\bm{\gamma}$. This is because the standard infill criteria must be adapted to cohesively incorporate both the surrogate model and the generative model simultaneously. Inspired by meta-learning principles [92, 93], we propose to reinterpret the infill criteria as an optimization loss for training the generative model. Without loss of generality, we define the generative model as $\bm{\gamma}(\cdot;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, parameterized by $\bm{\phi}$, which generates a new solution $\bm{x}_{t+1}$ from an existing input $\bm{x}$:
$\bm{x}_{t+1}=\bm{\gamma}\left(\bm{x};\operatorname{argmin}_{\bm{\phi}}\mathcal{L}(\bm{\phi})\right),$
$\text{where}\quad\mathcal{L}(\bm{\phi})=\alpha(\bm{\gamma}(\bm{x};\bm{\phi})).$ (9)
In this formulation, the infill criteria $\alpha(\cdot)$ explicitly defines the optimization objective used for training the parameters $\bm{\phi}$ of the generative model. Through this transformation, the task of minimizing the infill criteria is seamlessly integrated into the learning process of $\bm{\gamma}$, thereby allowing the model to internalize the required exploration-exploitation trade-offs during training.
Thus, by replacing the underlying objective function $f$ with its surrogate model $\tilde{f}$ in this generalized framework, we derive the resulting optimization loss for UCB/LCB as:
$\mathcal{L}_{\text{opt}}=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2},$ (10)
where $\mathbb{E}$ and $\mathbb{D}$ represent the mean and variance of the surrogate model’s output, respectively. This loss ties the generative model’s parameters directly to the surrogate’s predictions, facilitating a robust integration of the surrogate model into a generative modeling context. However, since the transformation of the infill criteria to the optimization loss is not guaranteed to be lossless, the inherent balance between exploration and exploitation may be compromised. Hence, we introduce an explicit variance optimization loss composed of UCB/LCB and Knowledge Gradient (KG) [94]:
$\mathcal{L}_{\text{opt}}$ $=\alpha_{\text{KG}}(\bm{p})$
$=\mathbb{E}\left[\mu^{*}_{t+1}-\mu^{*}_{t}\mid\bm{p}\right]+\lambda_{2}\mathbb{D}\left[\mu^{*}_{t+1}-\mu^{*}_{t}\mid\bm{p}\right]^{1/2}$
$=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2}+$
$\lambda_{2}\mathbb{C}\left[\tilde{f}_{t}(\bm{\gamma}^{\prime}\circ\bm{\gamma}(\bm{p})),\tilde{f}_{t}(\bm{q})\right]+\lambda_{2}\mathbb{C}\left[\tilde{f}_{t}(\bm{\gamma}\circ\bm{\gamma}^{\prime}(\bm{q})),\tilde{f}_{t}(\bm{p})\right],$ (11)
where $\mathbb{C}$ is the correlation function. Since the loss function is to be minimized, the existence of the correlation function requests that the parameter $\lambda_{2}$ is negative, which is not necessarily the case in Eq. 10. Supplementary Document B will elaborate the details of the justification of Eq. 11.
Unlike in conventional SAEAs, the surrogate model $\tilde{f}$ in our framework is not directly used to evaluate the quality of generated solutions. Instead, it serves only as a guidance mechanism during the training of the generative model. Furthermore, given the wide variety of infill criteria proposed in the literature, the infill criteria defined in Eq. 8 can be flexibly substituted with alternative strategies. This substitution naturally leads to different formulations of the optimization loss $\mathcal{L}_{\text{opt}}$, thereby allowing our framework to accommodate various exploration-exploitation trade-offs.
3) Training Procedure
The training procedure is summarized in Algorithm 2. It begins with training the surrogate model $\tilde{f}$ on the elite dataset $\mathcal{D}_{\text{SM}}$, after which the surrogate is frozen to ensure stable guidance throughout the optimization process. Subsequently, the paired generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are optimized using mini-batches sampled from the paired dataset $\mathcal{D}_{\text{GM}}$, which is constructed during the data preparation phase (Algorithm 1).
Algorithm 2 Model Training
1:Input: maximum training epochs $E$, optimizer opt, training datasets $\mathcal{D}_{\text{SM}}$ and $\mathcal{D}_{\text{GM}}$
2:Output: trained generative model $\bm{\gamma}$
[…]
4:for $e=1$ to $E$ do
5:  for each pair $((\bm{x}_{-},y_{-}),(\bm{x}_{+},y_{+}))$ in $\mathcal{D}_{\text{GM}}$ do
6:   $(\bm{p},y_{\bm{p}}),(\bm{q},y_{\bm{q}})\leftarrow(\bm{x}_{-},y_{-}),(\bm{x}_{+},y_{+})$
7:   $\mathcal{L}_{b}\leftarrow\mathcal{L}(\bm{\phi},\bm{\phi}^{\prime})$ where $\mathcal{L}$ is Eq. 1
8:   $\bm{\phi},\bm{\phi}^{\prime}\leftarrow\texttt{opt}(\bm{\phi},\nabla_{\bm{\phi}}\mathcal{L}_{b}),\texttt{opt}(\bm{\phi}^{\prime},\nabla_{\bm{\phi}^{\prime}}\mathcal{L}_{b})$
9:  end for
10:end for
11:Return: $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ where $\bm{\phi}^{*}\equiv\operatorname{argmin}_{\bm{\phi}}{\mathcal{L}}(\bm{\phi},\bm{\phi}^{\prime})$
In each training iteration (lines 4-10 of Algorithm 2), the model processes a paired sample $(\bm{x}_{-},y_{-})$ and $(\bm{x}_{+},y_{+})$ drawn from the inferior and superior subsets, respectively. The composite loss $\mathcal{L}$, as defined in Eq. 1, is evaluated and backpropagated to update the parameters of both generative models. Through this iterative process, the generative model $\bm{\gamma}$ progressively learns to map inferior solutions to superior ones, guided by the fitness landscape estimated by the surrogate model.
III-D Population Generation
Similar to the reproduction mechanism employed in traditional EAs, the population generation phase in EvoGO aims to create new candidate solutions based on the current population. However, instead of relying on handcrafted heuristic operators, EvoGO utilizes the learned generative model $\bm{\gamma}$ for solution generation.
Algorithm 3 Population Generation
1:Input: generative model $\bm{\gamma}$ with optimal parameter $\bm{\phi}^{*}$, latest population $\mathcal{D}_{t}$
2:Output: new population $\mathcal{D}_{t+1}$
[…]
4:for each $\left(\bm{x}_{i}^{t},y_{i}^{t}\right)\in\mathcal{D}_{t}$ do
5:  $\bm{x}_{i}^{t+1}\leftarrow\bm{\gamma}(\bm{x}_{i}^{t};\bm{\phi}^{*})$ // $\bm{\gamma}$” for trained generative model
6:  $y_{i}^{t+1}\leftarrow f(\bm{x}_{i}^{t+1})$ // Function evaluation
7:  $\mathcal{D}_{t+1}\leftarrow\mathcal{D}_{t+1}\cup\left\{\left(\bm{x}_{i}^{t+1},y_{i}^{t+1}\right)\right\}$
8:end for
As illustrated in Algorithm 3, this phase employs the trained generative model $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ to transform inferior solutions into superior ones, thereby generating new candidate solutions from the current population $\mathcal{D}_{t}$. Each generated solution $\bm{x}^{t+1}_{i}$ is subsequently evaluated using the true objective function $f$ to obtain its fitness value $y^{t+1}_{i}$, forming the next-generation population $\mathcal{D}_{t+1}$. Due to the inherently parallelizable nature of both the generative model and the objective function $f$, this entire population generation and evaluation step can be fully parallelized, which substantially improves computational efficiency.
IV Experiments
This section presents a comprehensive evaluation of EvoGO across diverse complex optimization tasks, including low- and high-dimensional numerical benchmarks, classical control problems, and complex high-dimensional robot control tasks. The experiments are rigorously designed to assess the effectiveness, efficiency, and scalability of our proposed data-driven evolutionary framework. Specifically, we aim to address the following research questions:
RQ1: How does EvoGO perform in terms of solution quality and convergence speed compared to representative baselines?
RQ2: What is the individual contribution of each major component in EvoGO to the overall performance?
RQ3: How does the optimization process of EvoGO evolve over time, and what can be observed from its search behavior through visual analysis?
To systematically address these research questions, we validate the proposed framework through three cohesive groups of experiments: (1) performance and efficiency comparisons across complex black-box optimization problems with varying dimensionalities; (2) ablation studies to analyze the impact of different components and hyperparameters; and (3) visualizations to reveal the behavioral characteristics of the learned optimization process. These studies collectively demonstrate the scalability, adaptability, and data-efficiency of EvoGO in solving complex optimization problems without reliance on heuristic operations.
IV-A Experimental Setup
Our empirical evaluation covers three categories of optimization challenges: (i) low- and high-dimensional numerical functions, (ii) low- and high-dimensional classical control tasks, and (iii) robotic control tasks. The classical control tasks Pushing and Rover are taken from [95], while Landing, Walker, and Ant are adapted from the Gym library [96]. The maximum number of fitness evaluations (FEs) is set to 1,000 for low-dimensional tasks and 10,000 for high-dimensional tasks.
EvoGO is compared against a comprehensive set of baseline methods, including: (i) Bayesian optimization algorithms, covering both sequential and batch variants, such as Hyperband [97], TuRBO [56], SCBO [98], BOGP (LogEI) [50], and the Tree structured Parzen Estimator (TPE) [99], which is widely used in AutoML; (ii) evolutionary strategies, represented by CMA-ES [11]; (iii) Nelder Mead [100], a classical heuristic method; (iv) SADE-SS [101] and SA-LSEO-LE [102], two state of the art surrogate assisted evolutionary algorithms.
All algorithms are evaluated over 20 independent runs with different random seeds on each benchmark. Additional details regarding the experimental environment, parameter configurations, and parallel execution settings are provided in Supplementary Document C, while ablation studies on key hyperparameters are included in Supplementary Document F.
(a) Rosenbrock (5)
(b) Levy (10D)
(c) Ackley (200D)
(d) Rastrigin (200D)
(e) Landing (12D)
(f) Pushing (14D)
(g) Walker (198D)
(h) Ant (404D)
(i) Plot legend
Figure 3: The convergence curves of EvoGO and the compared baselines in standard settings. The $y$-axes for numerical functions are in logarithmic scale, while those for other tasks are in linear scale since their true optimal is unknown. One-sigma error bars are adopted since two-sigma ones result in excessively large error bars for some algorithms. Please refer to Supplementary Document E for the complete figure.
IV-B Convergence Performance across Benchmarks
To address our first research question (RQ1), we evaluate the performance of EvoGO against a comprehensive set of representative baselines. This critical set of experiments is designed to rigorously assess the final solution quality and convergence speed across diverse tasks. Specifically, by testing on 25 benchmarks spanning numerical functions, classical control, and robot control, with dimensions ranging from 10D to 1000D, we aim to demonstrate the broad applicability and effectiveness of the method across various dimensionalities.
TABLE S.I: Statistical Results (Mean and Standard Deviation) of the Tested Algorithms on 1000-Dimensional Benchmark Problems. Each Algorithm Was Executed Ten Times, and the Best Results Are Highlighted in Bold.
Problem CMA-ES Nelder-Mead TuRBO SA_LSEO_LE EvoGO
Ackley 2.1192e+01 (2.66e-02) 2.1228e+01 (8.64e-03) 2.1236e+01 (2.80e-02) 2.0169e+01 (1.37e-01) 1.3652e+01 (1.10e-01)
Rosenbrock 1.4956e+09 (1.55e+08) 4.4097e+09 (1.73e+08) 5.9143e+08 (1.93e+07) 3.6834e+07 (1.32e+07) 5.8708e+06 (3.50e+05)
Rastrigin 9.2663e+05 (2.17e+04) 1.1251e+06 (1.73e+04) 4.2006e+05 (8.88e+03) 7.6479e+04 (7.78e+03) 5.4188e+04 (5.74e+02)
Levy 5.4255e+05 (3.66e+04) 5.8105e+05 (1.34e+04) 2.2472e+05 (7.57e+03) 1.6155e+05 (1.35e+04) 2.4058e+04 (8.77e+02)
The aggregated results, a representative subset of which is shown in Fig. 3 (with full details in Supplementary Document E), demonstrate that EvoGO consistently outperforms the baselines across most domains. A total of 25 benchmark tasks are evaluated, including four 1000-dimensional numerical problems listed in Table S.I. Due to the long runtime of high-dimensional cases, one representative comparison is provided for each algorithm. Overall, on low-dimensional problems, our method outperforms all baselines except the state-of-the-art SAEA algorithm SADE-SS, which can better exploit limited data in small-sample settings. However, for high-dimensional problems (e.g., 1000D), SADE-SS requires over one day of computation for a single run, making it impractical for large-scale applications. As shown in Table S.I, our method surpasses SA-LSEO-LE (state-of-the-art high-dimensional SAEA variants), indicating its superior capability to learn underlying regularities and distributions in complex black-box optimization tasks.
We further evaluate EvoGO on the Hopper environment provided by Brax [103], a high-dimensional robotic control benchmark that supports massively parallel simulation via GPU acceleration. As shown in Fig. 4, EvoGO achieves the highest final reward among all compared methods, including TPE, CMA-ES, and PPO [104], with respective population sizes of 1000, 1, 100, and 64. Notably, EvoGO attains comparable or even higher performance than PPO, despite the latter being an online reinforcement learning algorithm that continuously updates its policy through environment interaction. Furthermore, EvoGO exhibits substantial optimization efficiency: within a time budget of only 500 seconds, it surpasses all baselines, including those that continue training for over an hour. These results underscore the capability of EvoGO to deliver rapid and high-quality solutions in complex, GPU accelerated continuous control environments.
Figure 4: The rewards of TPE, CMA-ES, PPO, and EvoGO on the Hopper environment provided by Brax. To highlight the advantage in runtime of large population sizes, the rewards are plotted versus the wall-clock time, while the FEs are indicated by different markers. We limited the number of FEs and runtime to 10,000 and one hour respectively for all algorithms.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
Figure 5: The convergence curves of EvoGO and its upper-bound variant “Real Eval”, which directly utilizes real numerical functions instead of surrogate models.
IV-C Convergence Speed across Benchmarks
Further addressing our first research question (RQ1), this section focuses specifically on the convergence speed and efficiency of EvoGO. To this end, our experiments are designed from two angles: first, we measure the number of generations required for convergence on large-scale tasks, and second, we evaluate the wall-clock efficiency within a runtime-constrained environment.
First, as summarized in Table S.II, EvoGO converges within 10 generations across all large-scale benchmark tasks. In contrast, baseline methods often require tens to thousands of generations to achieve comparable performance. This property is crucial for scenarios with tight runtime budgets but abundant parallel resources, where large populations can be evaluated concurrently.
TABLE S.II: The number of generations to complete each benchmark by each method. Note: sequential methods are not listed here since they use exactly the same number of generations as FEs.
Benchmark TuRBO CMA-ES PSO Nelder-Mead EvoGO
[…]
Ant $4\sim25 1500\sim$1503 10
Second, the runtime constrained setting of Brax allows us to evaluate the wall clock efficiency of each method under the same budget of 10,000 function evaluations and one hour of runtime. As shown in Fig. 4, EvoGO rapidly attains a reward of 900 within the first 100 seconds, while other methods require substantially more time to reach comparable performance, and some never achieve it. In particular, EvoGO delivers up to $134\times$ higher reward per second than CMA ES and $50\times$ higher than TPE, reflecting its superior optimization throughput. These findings highlight the robust convergence rate of EvoGO when utilizing large population sizes and parallel evaluations.
IV-D Ablation Studies
To answer our second research question (RQ2), we conduct a comprehensive ablation study to dissect EvoGO and quantify the contribution of its key components. These experiments are designed to isolate the impact of elements such as the surrogate model, infill criteria, and generative architecture on overall performance.
(a) Ackley (200D)
(b) Rosenbrock (200D)
(c) Levy (200D)
(d) Walker (198D)
(e) Ant (404D)
(f) Plot Legend
Figure 6: The convergence curves of EvoGO and its ablations. “EvoGO” represents the original method; “Single Net” excludes the generative model $\bm{\gamma}$ and related loss terms; “Cycle GAN” adapts the Cycle GAN; “EvoGO-MLP” utilizes MLP for surrogate modeling; “EvoGO-LCB” employs the LCB infill criteria. Please refer to Supplementary Document F for the complete figure.
1) Under Perfect Surrogate Model
To examine the performance limit of EvoGO, we analyze its behavior when the surrogate model is assumed to be perfectly accurate. In this setting, the predicted objective values in the optimization loss are replaced with the ground-truth function evaluations, forming a variant referred to as Real Eval. This experiment isolates the effectiveness of the overall generative architecture and optimization strategy from any surrogate-induced error, thereby assessing the theoretical upper bound of the framework. As shown in Fig. 5, the Real Eval configuration achieves robust and consistent results across different numerical benchmarks, demonstrating the full potential of the proposed composite generative structure and loss formulation. These results indicate that EvoGO can achieve even higher performance when the surrogate model approaches perfect accuracy.
2) Under Failed Surrogate Model
To simulate the scenario where the surrogate model completely fails to provide reliable guidance, we evaluated two variants of EvoGO that remove or replace the surrogate component. The first variant, denoted as No Surrogate, eliminates the surrogate model and its corresponding optimization loss, relying solely on the generative mechanism for search guidance. The second variant, referred to as CycleGAN, replaces the surrogate-based optimization loss with a standard adversarial loss from the CycleGAN framework [80]. These two configurations effectively emulate conditions in which the surrogate model is either absent or incapable of meaningful prediction. As shown in Fig. 6, both variants exhibited noticeable performance degradation and increased instability, indicating that the surrogate model plays a crucial role in maintaining effective exploration-exploitation balance. Nevertheless, the generative component still preserved certain diversity and structural stability, largely due to the reconstruction constraint that prevents severe mode collapse. These observations confirm that while EvoGO can remain functional under surrogate failure, accurate surrogate feedback is vital for achieving its full performance potential.
3) Alternative Surrogate Model
To examine whether the proposed framework depends exclusively on GP surrogates, we conduct two complementary studies that evaluate both model substitution and infill criteria. These analyses reveal how EvoGO behaves when the GP-based probabilistic assumptions are replaced by alternative designs.
MLP as surrogate model. To evaluate the modularity of the surrogate component, we replaced the default GP with a Multi-Layer Perceptron (MLP) [105], forming the variant EvoGO-MLP. Since the MLP does not provide inherent uncertainty estimation, we removed the standard deviation and covariance terms from the optimization loss (Eq. 11). This configuration isolates the contribution of the probabilistic structure introduced by the GP. As shown in Fig. 6, the EvoGO-MLP variant showed slightly weaker performance on several numerical benchmarks, likely attributable to the absence of variance-guided exploration, while achieving strong results in complex scenarios such as the Ant benchmark. These results demonstrate that EvoGO remains effective even when the surrogate does not include explicit uncertainty modeling, highlighting the adaptability of the proposed framework.
infill criteria as replacement. We investigated the impact of the infill criteria by examining whether the proposed mechanism for balancing exploration and exploitation could be replaced by an alternative criterion. Specifically, we substituted the proposed optimization loss (Eq. 11) with the classical Lower/Upper Confidence Bound (LCB/UCB) criterion, thereby establishing the variant designated as EvoGO-LCB. The corresponding loss function is defined as
$\mathcal{L}_{\text{opt}}^{\text{LCB}}=\mathbb{E}\left[\tilde{f}_{t}(\bm{\gamma}(\bm{p}))\right]-y_{\bm{p}}+\lambda_{2}\mathbb{D}\left[\bm{\gamma}(\bm{p})\right]^{1/2},$ (12)
The derivation for the LCB infill criteria is provided in Supplementary Document B. As illustrated in Fig. 6, the “EvoGO-LCB” variant achieved competitive performance but generally performed worse than the original formulation. This observation indicates that, while the LCB/UCB-based infill criteria maintains reasonable performance, the tailored optimization loss employed in EvoGO provides a more stable exploration mechanism across diverse problem settings.
Generation 1
Generation 3
Generation 5
PH
PH
Fitness
Figure 7: Visual representation of vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$ on the Ackley function landscape. The vectors were generated by the trained generative model across different generations. They illustrate the transformation from input solutions $\bm{x}_{\text{in}}$ to the corresponding output solutions $\bm{x}_{\text{out}}$. The central star indicates the global optimum, while the dashed rectangles highlight regions covered by the solutions generated at different generations. For clarity, any landscape shifting and rotation have been omitted from this visualization.
4) Paired Architecture Analysis
We further analyze the structural importance of the paired generative architecture, which is a fundamental component of EvoGO. To this end, we design a “Single Net” variant that removes one generator and all consistency-enforcing mechanisms between $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$. Specifically, this variant eliminates the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5) and one branch of the similarity loss $\mathcal{L}_{\text{sim}}$ (Eq. 6) in Eq. 7, while omitting the standard deviation regularization term $\mathcal{L}_{\text{std}}$ in Eq. 11. This ablation assesses whether a single-generator design can sustain the stability and diversity achieved through paired modeling. As presented in Fig. 6, the “Single Net” configuration leads to a noticeable degradation in convergence stability and diversity maintenance, particularly under complex multimodal settings. These findings demonstrate that the paired architecture, reinforced by the reconstruction constraint, is essential for ensuring balanced training dynamics, preventing mode collapse, and enhancing the robustness of the overall generative process.
5) Extended Ablations
Additional ablation studies are detailed in Supplementary Document F. These studies examined variations in the neural network architectures and hyperparameter configurations, further elucidating the robustness and adaptability of EvoGO across various settings. Specifically, we explored the impact of different generative model layers and their width. We observed that the generative model’s core architecture has little effect on EvoGO’s performance, as shown in Fig. S4. Moreover, we also investigated the impact of the reconstruction loss scale $\lambda_{1}$, the data split factor $\eta$, and the sliding window factor $\epsilon$. These components were all found to have small or negligible impact on the performance of EvoGO, as shown in Figs. S5 to S7.
IV-E Search Dynamics
To answer our third research question (RQ3) and gain deeper insights into the search dynamics of EvoGO, we conducted a visualization study on a two-dimensional Ackley function with a population size of 100, as shown in Fig. 7. Specifically, we tracked the transitions from input solutions $\bm{x}_{\text{in}}$ to generated output solutions $\bm{x}_{\text{out}}=\bm{\gamma}(\bm{x}_{\text{in}})$, captured across various generations from the trained generative model $\bm{\gamma}$. Arrows in the visualization represent vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$, and their colors reflect their norm-2 lengths, as described in the corresponding plot legends. This visual representation demonstrates EvoGO’s capacity to precisely estimate both the direction $(\bm{x}_{\text{out}}-\bm{x}_{\text{in}})/\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ and the magnitude $\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ of steps toward potential optima. This capability not only enhances optimization efficiency but also significantly mitigates mode collapse. Furthermore, the learning progress of $\bm{\gamma}$ is illustrated through the grayscale contours of the optimization landscapes captured by the solutions at different generations. Throughout the search process, $\bm{\gamma}$ adaptively hones in on specific regions, thereby demonstrating a refined focus that enhances its search efficacy. This adaptiveness suggests EvoGO’s potential for achieving superlinear convergence given ample data, which is accentuated by the dashed rectangles that delineate the boundaries of solutions generated at different generations.
V Conclusion
This paper presented Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework. This framework was specifically designed to address three critical limitations inherent in existing generative model-based evolutionary algorithms: reliance on handcrafted heuristics, objective misalignment between generative models and evolutionary goals, and underutilization of available data. By eliminating traditional search operators and adopting globally consistent training objectives, EvoGO enables robust end-to-end learning for optimization, offering enhanced generality, objective alignment, and data efficiency.
The proposed framework is underpinned by a paired generative architecture that integrates forward and inverse models to produce high-fidelity candidate solutions, thereby effectively replacing conventional crossover and mutation operators. To enhance training diversity under limited evaluation budgets, we employed a pairwise data construction strategy. Furthermore, a composite loss function integrates reconstruction, distributional alignment, and directional guidance to ensure efficient model convergence toward promising objective regions.
Extensive experiments conducted across numerical, control, and robotic tasks demonstrated the superior performance of EvoGO when compared against state-of-the-art baselines. Specifically, EvoGO consistently achieved convergence within 10 generations. This performance, combined with GPU-parallelized inference, delivered up to a $134\times$ speedup on computationally complex optimization tasks.
These results underscore the substantial potential of fully data-driven evolutionary frameworks to generalize effectively across diverse domains without requiring manual heuristic design. Future work will initially focus on enhancing robustness under extremely low data regimes and providing theoretical insights into the framework’s convergence behavior and optimization dynamics. In addition, future extensions will explore offline pre-training and cross-task transfer to further improve data efficiency in settings where online evaluation is limited.